### v1

# SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models

Abstract
The rapid advancement in large language models (LLMs) comes with a significant increase in their parameter size, presenting challenges for adaptation and fine-tuning. Parameter-efficient fine-tuning (PEFT) methods are widely used to adapt LLMs for downstream tasks efficiently. In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method. We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) on weights and discuss SORSA’s superiority in minimizing the deviation from the pre-trained weight. Each SORSA layer consists of two main parts: trainable principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These parts are initialized by performing SVD on pre-trained weights. Moreover, we implement an orthonormal regularizer and analyze its importance by performing gradient analysis. The analysis shows that the regularizer could effectively transfer the scaling information into $\Sigma_{p}$, which ensures the parameter updating of SORSA layers is evenly and minimized on $U_{p}$ and $V^{\top}_{p}$. SORSA layers could be merged during inference, thus eliminating inference latency. After all, SORSA shows a faster convergence speed than PiSSA and LoRA in our experiments. On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%). On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%) We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance. The code is available at https://github.com/Gunale0926/SORSA.
Deep Learning, LLM, PEFT, LoRA, SVD
Figure 1: Architecture of a SORSA layer. We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
1 Introduction
Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks. For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
However, as the number of parameters in LLMs rapidly increases, full parameter fine-tuning becomes increasingly inefficient. For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology. Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB, thus rendering it unfeasible on a single GPU with current technology.
To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed. These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning. For instance, tuning Llama 2 7B in Float32 by LoRA with a rank of 128 only takes approximately 60GB VRAM, which allows training in 1 $\times$ NVIDIA A100-80GB or 3 $\times$ NVIDIA RTX 4090-24GB.
Among those PEFT methods, LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023; Meng et al., 2024; Liu et al., 2024; Dettmers et al., 2024) had become increasingly popular due to their: 1. low training VRAM requirement 2. no inference latency 3. versatility in different neuron network architectures.
In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA. A SORSA layer has two main parts: principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$. These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight. Residual singular values and vectors will be merged into one matrix and frozen while training. We only train principle singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$. The architecture of a SORSA layer is illustrated in Figure 1.
Furthermore, we analyze the pattern of variance of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
2 Related Works
Parameter-efficient fine-tuning (PEFT) methods have been developed to address the inefficiency of full parameter fine-tuning for large language models. These methods focus on adapting the model for downstream tasks while updating only a few parameters and keeping most of the model’s weights frozen. This approach significantly reduces the memory and computational requirements during training, especially VRAM.
Adapter-based PEFT is the first type of PEFT, which was initially designed by (Houlsby et al., 2019). It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters. Its variants, e.g., (Lin et al., 2020), reduce the number of adapter layers per block, and (He et al., 2022) focus on adding adapter modules parallel to existing layers. However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in (Lester et al., 2021). There are several variants of this work, including (Liu et al., 2022; Razdaibiedina et al., 2023). However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
LoRA (Hu et al., 2021) and its variants are the most popular type of PEFT methods. This type of PEFT is popular for its on-par or better performance than full parameter fine-tuning without introducing any inference latency. LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices. Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating; DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm. The results show that DoRA has a better learning capacity than LoRA. However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA; PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}$ into frozen $W_{pri}=AB$ and $W_{res}=W_{0}-W_{pri}$ based on singular values, and only updating the most important part $W_{pri}$, which result a faster convergence speed and better fitting than LoRA. SORSA has a similar architecture to PiSSA, which both conducts singular value decomposition (SVD) and replaces pre-trained weights with residual singular weights. SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
Other methods. There are also a few PEFT methods with unique techniques. For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation. Despite its efficiency, GaLore’s complex implementation, involving Singular Value Decomposition (SVD) on gradients, adds computational complexity and may face scalability issues for very large models. LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance. However, LISA’s selective fine-tuning requires careful resource allocation and introduces additional training complexity, complicating the pipeline and necessitating more sophisticated monitoring mechanisms.
3 Singular Values and Vectors Analysis
3.1 Singular Value Decomposition
The geometric meaning of SVD can be summarized as follows: for every linear mapping $W:K^{m}\to K^{n}$, the singular value decomposition finds a set of orthonormal bases in both the original space and the image space such that $W$ maps the $i$-th basis vector of $K^{m}$ to a non-negative multiple of the $i$-th basis vector of $K^{n}$, and maps the remaining basis vectors of $K^{m}$ to the zero vector. In other words, the matrix $W$ can be represented as $\Sigma$ on these selected bases, where $\Sigma$ is a diagonal matrix with all non-negative diagonal elements.
According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\Sigma V^{\top}$. Here, $U\in\mathbb{R}^{m\times k}$ are left singular vectors and have orthonormal columns, $V^{\top}\in\mathbb{R}^{k\times n}$ are right singular vectors and have orthonormal rows, and $\Sigma\in\mathbb{R}^{k\times k}$ is a diagonal matrix, where diagonal values are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order.
3.2 Analysis Method
The study of DoRA (Liu et al., 2024) introduces an analysis method that focuses on the deviation of magnitude and direction ($\Delta M,\Delta D$) during training of full parameter fine-tuning and LoRA (Hu et al., 2021). They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference. Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices. Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m,n}$ and tuned weight $W_{t}\in\mathbb{R}^{m,n}$, which $t$ denotes the training step, could be defined as follows:
$\Delta\Sigma_{t}=\frac{\Sigma_{i=1}^{k}\left|\sigma_{t}^{i}-\sigma_{0}^{i}\right|}{k}$ (1)
Here, $\Delta\Sigma_{t}$ represents singular value variants between $W_{0}$ and $W_{t}$ at training step $t$. $\sigma^{i}_{t}$ denotes the $i$-th element in diagonal of $\Sigma_{t}$, where $\Sigma_{t}$ is decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$;
$\Delta U_{j}=\left|\sum_{i=1}^{m}\left(U_{t}\odot U_{0}\right)_{ij}\right|$ (2)
$\Delta V^{\top}_{i}=\left|\sum_{j=1}^{n}\left(V_{t}^{\top}\odot V_{0}^{\top}\right)_{ij}\right|$ (3)
$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{i}+\Delta V^{\top}_{i})$ (4)
Here, $\Delta D_{t}\in(0,1)$ represents variants of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$. $\Delta U\in\mathbb{R}^{k}$, $\Delta V^{\top}\in\mathbb{R}^{k}$. $\odot$ denotes element-wise product, and $\left|\cdot\right|$ denotes element-wise absolute value. $U_{t}$ and $V_{t}$ are decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$.
We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024). We test partial fine-tuning, LoRA, and SORSA (with regularizer with $\gamma=0.1$ and without regularizer). See Appendix A.1 for training details of the analysis.
3.3 Analysis Result
This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2. The four methods obtained similar final loss ($\pm 10\%$).
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 2: $\Delta D$ and $\Delta\Sigma$ of each trainable parameters during training steps. Numbers in the plot represent layer of the weight. Dots represent mean $\Delta D$ and $\Delta\Sigma$ at specific step.
The partial fine-tuning and LoRA methods analysis show that both methods exhibit significant adjustments in significant vectors $\Delta D$. This substantial alteration disrupts the characteristics of the pre-trained matrix and is likely to affect the model’s generalization ability. Moreover, the updates of all parameters in both partial fine-tuning and LoRA methods show a parallel updating pattern across weights in different layers, which emphasizes a restriction of these methods and may lead to a potentially disruptive modification of the generalization ability.
SORSA with regularizer shows $\Delta D$ and $\Delta\Sigma$ are in a much smaller range than other analysis methods. Moreover, different matrices demonstrated unrelated updating patterns, compared to all parallel-like other three methods, indicating that the updates in the SORSA are less constrained and potentially converge faster. In contrast, SORSA without the orthonormal regularizer exhibits a greater $\Delta D$ and $\Delta\Sigma$ alteration at each training step compared to SORSA with the regularizer. SORSA without regularizer also shows a linear-like updating pattern, similar to LoRA and partial FT. The result verifies the importance of regularizer in maintaining stability and minimizing the deviation from pre-trained matrices during the training process. These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively. This property allows SORSA layers to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
4 Our Method
4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
According to the definition of SVD in Section 3.1, given a rank $r$ where $r\ll k$, we could perform the low-rank approximation by selecting the first $r$ items on the diagonal of $\Sigma$, which is the first $r$ most significant singular values, and also select the first $r$ columns of $U$ and first $r$ rows of $V^{\top}$, which correspond to the selected singular values. By performing SVD low-rank approximation, we could get a low-rank matrix that preserves the largest significant values and vectors, containing the matrix’s most significant data.
Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principle weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
$W_{0}=W_{p}+W_{r}$ (5)
$W_{p}=U_{[:,:r]}\Sigma_{[:r,:r]}V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (6)
$W_{r}=U_{[:,r:]}\Sigma_{[r:,r:]}V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (7)
Here, $U$ represents the left singular vectors, $\Sigma$ represents the diagonal matrix of singular values, and $V^{\top}$ represents the right singular vectors. We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix. We rewrite $U_{[:,:r]}$, $\Sigma_{[:r,:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $\Sigma_{[r:,r:]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$ correspondingly.
The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024). Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices. SORSA is defined by Eq (8), which is initially equivalent to the pre-trained weight $W_{0}$.
During training, $W_{r}$ remains frozen, and only $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ are updated. Since $\Sigma_{p}$ is a diagonal matrix, only the diagonal elements are updated, while the rest of the matrix remains zero.
$\text{SORSA}(x):=W_{r}x+W_{p}x$ (8)
We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 4.2.
$\mathcal{L}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V_{p}V^{\top}_{p}-I\|_{F}$ (9)
Where $\mathcal{L}_{reg}$ is the orthonormal regularizer loss, the $U_{p}$ and $V^{\top}_{p}$ are each orthonormal vectors in columns and rows, respectively, after initialization due to SVD’s property. The regularizer could keep their orthonormality during training.
Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA layer at training step $t$ could be expressed as:
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}\mathcal{L}_{train}-\gamma_{t}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (10)
At training step $t$, $\nabla_{W_{t}}\mathcal{L}_{train}$ denotes the gradient of $W_{t}$ respect to $\mathcal{L}_{train}$, and $\nabla_{W_{t}}\mathcal{L}_{reg}$ denotes the gradient of $W_{t}$ respect to the orthonormal regularizer loss $\mathcal{L}_{reg}$. $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss, respectively.
We update the SORSA as Eq (11) for implementation simplicity.
$W_{p,t+1}=$ $W_{p,t}-\eta_{t}\left(\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}\right)$ (11)
$\eta_{d}$ is the learning rate which is given to the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
4.2 Gradient Analysis
In this section, we analyze the impact of an orthonormal regularization on the updates of the singular vector matrix $U_{p}$ and $V^{\top}_{p}$ during training and the importance of keeping $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ separately. We also analyze the scaling factor for the learning rate of $\Sigma_{p}$
Given $U_{p,t-1}\in\mathbb{R}^{m\times r}$ is an orthonormal matrix, which corresponds to the initial state of $U_{p}$ due to the property of SVD. Here, we denote the gradient without regularizer for matrix $W$ at step $t$ is by $\nabla_{W_{t}}$, and the gradient with regularizer by $\nabla^{\prime}_{W_{t}}$:
$\nabla_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (12)
$\nabla^{\prime}_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}+\gamma\nabla_{U_{p,t-1}}\mathcal{L}_{reg}$ (13)
Where $\nabla_{U_{p,t-1}}\mathcal{L}_{reg}=\frac{\partial\mathcal{L}_{reg}}{\partial U_{p,t-1}}=0$ due to the orthonormality of $U_{p,t-1}$; $\nabla_{U_{p,t-1}}\mathcal{L}_{train}=\frac{\partial\mathcal{L}_{train}}{\partial U_{p,t-1}}$.
Thus, we could calculate the weight at step $t$ without regularizer $U_{t}$, and with regularizer $U_{t}^{\prime}$:
$U_{p,t}^{\prime}=U_{p,t}=U_{p,t-1}-\eta\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (14)
Based on their weight on step $t$, we could calculate $W_{t+1}$ and $W^{\prime}_{t+1}$. Here, we only focus on the gradient of $U_{p}$ and $V^{\top}_{p}$, thus ignoring the updating of $\Sigma_{p,t}$ in this step.
$U_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}$ (15)
$V^{\top}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}$ (16)
$W_{t+1}=U_{t+1}\Sigma_{p,t+1}V^{\top}_{t+1}$ (17)
$U^{\prime}_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}$ (18)
$V^{\top\prime}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}$ (19)
$W^{\prime}_{t+1}=U^{\prime}_{t+1}\Sigma_{p,t+1}V^{\top\prime}_{t+1}$ (20)
Under optimal circumstances, $U^{\prime}_{p,t+1}$ will be an orthonormal matrix. Therefore, $U^{\prime}_{p,t+1}$ can be interpreted as pure rotation, which does not change the norm of the input matrix. By enforcing the orthonormal property of $U_{p}$ and $V^{\top}_{p}$ at each step, their regularized gradients $\nabla_{U_{p}}\mathcal{L}_{reg}$ and $\nabla_{V^{\top}_{p}}\mathcal{L}_{reg}$ will ensuring that the scaling information is not propagated into $U_{p,t}$ and $V^{\top}_{p,t}$. Consequently, only $\Sigma_{p}$ retains all scaling information.
To prove this statement, we could calculate the difference between $W^{\prime}_{t+1}$ and $W_{t+1}$:
$\begin{split}W^{\prime}_{t+1}-W_{t+1}&=-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}U_{p,t}\Sigma_{p,t+1}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{train}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\\ &\quad-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}V^{\top}_{p,t}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}\Sigma_{p,t+1}\\ &\quad+\gamma^{2}\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\end{split}$ (21)
Here, we could find that each term has the $\Sigma_{p,t+1}$. This shows that by optimizing respected to regularizer loss at step $t+1$ and optimizing $\Sigma_{p,t+1}$ respected to train loss on step $t+2$, we could efficiently increase the orthonormality of $U_{p}$ and $V^{\top}_{p}$ while have negligible side-effect on convergence. In short, adding regularizer could transfer scaling information from $U_{p}$ and $V^{\top}_{p}$, which come from optimizing for $\mathcal{L}_{train}$, into $\Sigma_{p}$ allows the matrix scaling information updates more holistically, ultimately presenting a better generalization. In Figure 3, we illustrate the effect of $\gamma$ on orthonormality, training loss, and gradient norm. The figures present almost no difference in loss and gradient norm in different $\gamma$ values, which indicates the strong robustness of training of SORSA.
Figure 3: Orthonormality, loss and gradient norm variation of SORSA using different $\gamma$ on MetaMathQA training of Mistral 7B v0.1.
Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$. By ensuring the scaling information only be updated into $\Sigma_{p}$, we could also ensure that the parameter distribution of $W_{p}$ is evenly distributed.
Introducing the orthonormal regularizer into the training process promotes a more stable update trajectory and a more evenly distributed weight by actively countering deviations from orthonormality. This regulation stabilizes the updates and enhances the model’s ability to maintain crucial properties of the singular vectors $U_{p}$ and $V^{\top}_{p}$, thereby contributing to improved generalization and effectiveness in learning. This analysis underlines the utility of orthonormal regularization in complex model training scenarios, particularly in fine-tuning pre-trained models where preserving learned features is crucial.
5 Empirical Experiments
We conduct comparative experiments on different NLP tasks, including natural language generation (NLG) between SORSA, PiSSA (Meng et al., 2024), LoRA (Hu et al., 2021), and full parameter fine-tuning.
We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023). We trained the models with SORSA applied and conducted GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) tests. The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024). See Appendix A.2 for more details and hyperparameters of the training. We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from the PiSSA paper (Meng et al., 2024). For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80G) GPU. See Table 1 for the results and Figure 4 for the loss and gradient norm comparison. See Figure 5 for the variation in orthonormality of SORSA.
Model Method Trainable Parameters MATH GSM-8K
Llama 2 7B Full FT 6738M 7.22 49.05
LoRA 320M 5.50 42.30
PiSSA 320M 7.44 53.07
SORSA 320M 10.36 56.03
Mistral 7B Full FT 7242M 18.60 67.02
LoRA 168M 19.68 67.70
PiSSA 168M 21.54 72.86
SORSA 168M 21.86 73.09
Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from the paper of PiSSA (Meng et al., 2024).
Refer to caption Refer to caption
Figure 4: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B.
Figure 5: Variation of orthonormality on MetaMathQA training of Llama 2 7B and Mistral 7B v0.1.
The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT. On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both tasks compared to the other three methods.
In this section, due to the limitation of computing resources, we train and benchmark a small number of tasks.
6 Conclusion
In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks. SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principle and residual components, only training the principle singular values and vectors while freezing the residuals. We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
Our experiments demonstrated that SORSA outperforms existing PEFT methods, such as LoRA and PiSSA, in both convergence speed and accuracy on the NLG tasks. Specifically, Llama 2 7B, tuned with SORSA, achieved significant improvements in the GSM-8K and MATH benchmarks, highlighting the effectiveness of our approach.
We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training. This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment. We also show the significance of the orthonormal regularizer through analysis. We also introduce the gradient analysis of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
SORSA retains the advantages of LoRA and variants, including low training VRAM requirements, no inference latency, and versatility across different neural network architectures. By offering a more efficient fine-tuning mechanism, SORSA presents a promising direction for future research and application in the field of LLMs.
Overall, SORSA gives a new perspective on parameter-efficient fine-tuning, showcasing exceptional efficiency and robust performance. It not only outperforms existing methods like LoRA and PiSSA in several downstream tasks but also maintains the practical benefits of low VRAM requirements, no inference latency, and ease of implementation. This innovative approach offers a promising direction for future research and practical applications in adapting pre-trained models, making it a pivotal development in the field.
7 Future Work
While SORSA demonstrates substantial improvements over existing PEFT methods, there are several avenues for future work to enhance its capabilities further and extend its applicability:
  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different transformer-based LLMs. This includes replicating other tests from studies like PiSSA (Meng et al., 2024) and performing evaluations such as HumanEval (Chen et al., 2021) and MT-Bench (Zheng et al., 2023), in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2021a) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
  • Application to Other Fields: Test SORSA on different fields beyond NLP. Such as conducting experiments on computer vision (CV) models like DDPM (Ho et al., 2020), DDIM (Song et al., 2020), and ViT (Dosovitskiy et al., 2021). These experiments could expand SORSA’s applicability to a broader range of deep-learning tasks.
  • Quantized SORSA (QSORSA): Investigate the application of quantization techniques to SORSA introduced in QLoRA (Dettmers et al., 2024). This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources. By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
By pursuing these directions, future work can build on the foundation laid by SORSA, pushing the boundaries of PEFT and enhancing the adaptability and performance of large language models across a wide range of applications. This will enable the development of more versatile downstream models, potentially expanding the impact of Machine Learning into more fields and becoming an integral part of people’s everyday lives.
8 Impact Statement
In this paper, we introduce an innovative PEFT method in the area of Machine Learning. Our approach significantly streamlines the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability. As we push the boundaries of what is possible with machine learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
8.1 Environmental Impact
SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B model than full parameter fine-tuning. This significant reduction in hardware resource requirements translates to approximately 50% less energy consumption compared to full parameter fine-tuning methods. By enhancing efficiency, our approach contributes to reducing the environmental footprint of machine learning operations.
8.2 Ethical Concerns
The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models. As demonstrated in (Lermen & Rogers-Smith, 2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content. The ease and affordability of such methods underscore the vulnerability of safety protocols. It is imperative to develop robust safeguards that keep pace with technological advancements, ensuring that efficiency gains in model tuning do not compromise the ethical use of AI.
9 Acknowledgment
We would like to extend our gratitude to Prof. Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing guidance to the research.
Appendix A Training Details
A.1 Analysis
For the singular values and vectors analysis in section 3, we apply partial fine-tuning, LoRA SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, using the same training data and optimizer as A.2. In the analysis, LoRA and SORSA are only applied to q_proj and v_proj matrices, respectively. FT is only trainable on q_proj and v_proj. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 2 for hyperparameters.
Training
Method FT LoRA SORSA (w/o reg) SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
Learning Rate 2e-5 2e-5 2e-5 3e-5
$\gamma$ 0 5e-4
Rank 128 128 128
Table 2: Hyperparameters for the analysis
A.2 NLG Experiments
For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training. SORSA layers are applied on all linear matrices in every layer. We only calculate the loss on the response part. The models are trained with BF16 mix precision. See Table 3 for hyperparameters.
Training
Model Llama 2 7B Mistral 7B v0.1
Method SORSA
Precision TF32 + BF16 Mix-Precision
Epoch 1
Batch Size 128
Max Length 512
Weight Decay 0.00
Warm-up Ratio 0.03
Learning Rate 3e-5
$\gamma$ 4e-4
Rank 128 64
Inference
Precision BF16
Top-P 1.0
GSM-8K Max Length 1024
MATH Max Length 2048
Table 3: Hyperparameters of training on models for GSM-8K and MATH

### v2

@@
 In this paper, we propose Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA, a novel PEFT method.
-We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) on weights and discuss SORSA’s superiority in minimizing the deviation from the pre-trained weight.
-Each SORSA layer consists of two main parts: trainable principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
+We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect.
+Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
 These parts are initialized by performing SVD on pre-trained weights.
-Moreover, we implement an orthonormal regularizer and analyze its importance by performing gradient analysis.
-The analysis shows that the regularizer could effectively transfer the scaling information into $\Sigma_{p}$, which ensures the parameter updating of SORSA layers is evenly and minimized on $U_{p}$ and $V^{\top}_{p}$.
-SORSA layers could be merged during inference, thus eliminating inference latency.
-After all, SORSA shows a faster convergence speed than PiSSA and LoRA in our experiments.
+Moreover, we implement and analyze an orthonormal regularizer, which could effectively transfer the scaling information into $\Sigma_{p}$ and ultimately allows the training process to be more efficient.
+SORSA adapters could be merged during inference, thus eliminating any inference latency.
+After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments.
 On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%).
-On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%) We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance.
+On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%).
+We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance.
 The code is available at https://github.com/Gunale0926/SORSA.
 Deep Learning, LLM, PEFT, LoRA, SVD
-Figure 1: Architecture of a SORSA layer.
+Figure 1: Architecture of a SORSA adapter.
 We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
@@
 In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA.
-A SORSA layer has two main parts: principle singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
+A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
 These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight.
 Residual singular values and vectors will be merged into one matrix and frozen while training.
-We only train principle singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$.
-The architecture of a SORSA layer is illustrated in Figure 1.
-Furthermore, we analyze the pattern of variance of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
+We only train principal singular values and vectors with an orthonormal regularizer implemented to keep the orthonormality of $U_{p}$ and $V_{p}^{\top}$.
+The architecture of a SORSA adapter is illustrated in Figure 1.
+Furthermore, we analyze the pattern of variation of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
 In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
@@
 LoRA could be represented by equation $W=W_{0}+BA$, where $W_{0}\in\mathbb{R}^{m\times n}$ is the pre-trained weight, $A\in\mathbb{R}^{m\times r}$, using Gaussian initialization, and $B\in\mathbb{R}^{r\times n}$, using zero initialization, are low-rank matrices.
-Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating; DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm.
+Its variant, for example, AdaLoRA (Zhang et al., 2023) introduces an SVD decomposition and pruning for least significant singular values for more efficient parameter updating.
+DoRA (Liu et al., 2024) proposed a novel way to decompose weight into direction and magnitude by $W=\underline{m}\frac{W_{0}+\underline{BA}}{\|W_{0}+\underline{BA}\|_{c}}$, where $\underline{m}$ is initialized by $\underline{m}=\|W_{0}+\underline{BA}\|_{c}$, $\|\cdot\|_{c}$ denotes column-wise norm.
 The results show that DoRA has a better learning capacity than LoRA.
-However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA; PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}$ into frozen $W_{pri}=AB$ and $W_{res}=W_{0}-W_{pri}$ based on singular values, and only updating the most important part $W_{pri}$, which result a faster convergence speed and better fitting than LoRA.
-SORSA has a similar architecture to PiSSA, which both conducts singular value decomposition (SVD) and replaces pre-trained weights with residual singular weights.
+However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA.
+PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=BA$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen.
+PiSSA results a faster convergence speed and better fitting compared to LoRA.
+SORSA has a similar architecture to PiSSA, which both conducts SVD and replaces pre-trained weights with residual singular weights.
 SORSA inherits LoRA and its variants’ benefits, including low training VRAM requirement, no inference burden, and versatility in different architectures.
 Other methods.
-There are also a few PEFT methods with unique techniques.
+There are also a few efficient adapting methods with unique techniques.
 For example, GaLore (Zhao et al., 2024) is a memory-efficient PEFT method that reduces VRAM usage by leveraging gradient accumulation and low-rank approximation.
-Despite its efficiency, GaLore’s complex implementation, involving Singular Value Decomposition (SVD) on gradients, adds computational complexity and may face scalability issues for very large models.
-LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance.
-However, LISA’s selective fine-tuning requires careful resource allocation and introduces additional training complexity, complicating the pipeline and necessitating more sophisticated monitoring mechanisms.
+Despite its efficiency, GaLore’s complex implementation, involving perform SVD on gradients, adds computational complexity and may face scalability issues for very large models.
+LISA (Pan et al., 2024) uses a layer-wise importance sampling approach, prioritizing layers that significantly impact model performance, and selectively fine-tune essential parameters.
 3 Singular Values and Vectors Analysis
@@
 They discovered that the distinction between full parameter fine-tuning and LoRA likely affects their learning ability difference.
-Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices.
+Inspired by their methods, we propose a novel method that analyzes the correlation between the deviation of singular values ($\Delta\Sigma$) and singular vectors ($\Delta D$) from pre-trained matrices during updating.
 Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
@@
 $\sigma^{i}_{t}$ denotes the $i$-th element in diagonal of $\Sigma_{t}$, where $\Sigma_{t}$ is decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$;
-$\Delta U_{j}=\left|\sum_{i=1}^{m}\left(U_{t}\odot U_{0}\right)_{ij}\right|$ (2)
-$\Delta V^{\top}_{i}=\left|\sum_{j=1}^{n}\left(V_{t}^{\top}\odot V_{0}^{\top}\right)_{ij}\right|$ (3)
-$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{i}+\Delta V^{\top}_{i})$ (4)
-Here, $\Delta D_{t}\in(0,1)$ represents variants of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$.
-$\Delta U\in\mathbb{R}^{k}$, $\Delta V^{\top}\in\mathbb{R}^{k}$.
-$\odot$ denotes element-wise product, and $\left|\cdot\right|$ denotes element-wise absolute value.
-$U_{t}$ and $V_{t}$ are decomposed from $W_{t}$ by performing SVD, $k=\min(m,n)$.
+$\Delta U_{t,j}=\left|\langle\mathbf{u}_{t}^{j},\mathbf{u}_{0}^{j}\rangle\right|$ (2)
+$\Delta V^{\top}_{t,i}=\left|\langle\mathbf{v}_{t}^{i},\mathbf{v}_{0}^{i}\rangle\right|$ (3)
+$\Delta D_{t}=1-\frac{1}{2k}\sum_{i=0}^{k}(\Delta U_{t,i}+\Delta V^{\top}_{t,i})$ (4)
+Here, $k=\min(m,n)$; $\mathbf{u}_{t}^{j}$ denotes the $j$-th column vector of matrix $U_{t}$, and $\mathbf{v}_{t}^{i}$ denotes the $i$-th row vector of matrix $V_{t}^{\top}$; $\Delta D_{t}\in(0,1)$ represents variation of singular vectors between $W_{0}$ and $W_{t}$ at training step $t$; $U_{t}$ and $V_{t}^{\top}$ are decomposed from $W_{t}$ by performing SVD.
 We adopt the analysis on Llama 2 7B (Touvron et al., 2023) using the first 100K data of MetaMathQA (Yu et al., 2024).
-We test partial fine-tuning, LoRA, and SORSA (with regularizer with $\gamma=0.1$ and without regularizer).
+We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer).
 See Appendix A.1 for training details of the analysis.
@@
 These updating patterns indicate that SORSA maintains the characteristics of the pre-trained matrix better, thus potentially preserving the model’s generalization ability more effectively.
-This property allows SORSA layers to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
+This property allows SORSA adapters to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
 4 Our Method
@@
 By performing SVD low-rank approximation, we could get a low-rank matrix that preserves the largest significant values and vectors, containing the matrix’s most significant data.
-Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principle weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
+Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principal weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
 $W_{0}=W_{p}+W_{r}$ (5)
@@
 The regularizer could keep their orthonormality during training.
-Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA layer at training step $t$ could be expressed as:
+Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
 $W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}\mathcal{L}_{train}-\gamma_{t}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (10)
@@
 4.2 Gradient Analysis
-In this section, we analyze the impact of an orthonormal regularization on the updates of the singular vector matrix $U_{p}$ and $V^{\top}_{p}$ during training and the importance of keeping $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ separately.
-We also analyze the scaling factor for the learning rate of $\Sigma_{p}$
-Given $U_{p,t-1}\in\mathbb{R}^{m\times r}$ is an orthonormal matrix, which corresponds to the initial state of $U_{p}$ due to the property of SVD.
-Here, we denote the gradient without regularizer for matrix $W$ at step $t$ is by $\nabla_{W_{t}}$, and the gradient with regularizer by $\nabla^{\prime}_{W_{t}}$:
-$\nabla_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (12)
-$\nabla^{\prime}_{U_{p,t-1}}=\nabla_{U_{p,t-1}}\mathcal{L}_{train}+\gamma\nabla_{U_{p,t-1}}\mathcal{L}_{reg}$ (13)
-Where $\nabla_{U_{p,t-1}}\mathcal{L}_{reg}=\frac{\partial\mathcal{L}_{reg}}{\partial U_{p,t-1}}=0$ due to the orthonormality of $U_{p,t-1}$; $\nabla_{U_{p,t-1}}\mathcal{L}_{train}=\frac{\partial\mathcal{L}_{train}}{\partial U_{p,t-1}}$.
-Thus, we could calculate the weight at step $t$ without regularizer $U_{t}$, and with regularizer $U_{t}^{\prime}$:
-$U_{p,t}^{\prime}=U_{p,t}=U_{p,t-1}-\eta\nabla_{U_{p,t-1}}\mathcal{L}_{train}$ (14)
-Based on their weight on step $t$, we could calculate $W_{t+1}$ and $W^{\prime}_{t+1}$.
-Here, we only focus on the gradient of $U_{p}$ and $V^{\top}_{p}$, thus ignoring the updating of $\Sigma_{p,t}$ in this step.
-$U_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}$ (15)
-$V^{\top}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}$ (16)
-$W_{t+1}=U_{t+1}\Sigma_{p,t+1}V^{\top}_{t+1}$ (17)
-$U^{\prime}_{t+1}$ $=U_{p,t}-\eta\nabla_{U_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}$ (18)
-$V^{\top\prime}_{t+1}$ $=V^{\top}_{p,t}-\eta\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}$ (19)
-$W^{\prime}_{t+1}=U^{\prime}_{t+1}\Sigma_{p,t+1}V^{\top\prime}_{t+1}$ (20)
-Under optimal circumstances, $U^{\prime}_{p,t+1}$ will be an orthonormal matrix.
-Therefore, $U^{\prime}_{p,t+1}$ can be interpreted as pure rotation, which does not change the norm of the input matrix.
-By enforcing the orthonormal property of $U_{p}$ and $V^{\top}_{p}$ at each step, their regularized gradients $\nabla_{U_{p}}\mathcal{L}_{reg}$ and $\nabla_{V^{\top}_{p}}\mathcal{L}_{reg}$ will ensuring that the scaling information is not propagated into $U_{p,t}$ and $V^{\top}_{p,t}$.
-Consequently, only $\Sigma_{p}$ retains all scaling information.
-To prove this statement, we could calculate the difference between $W^{\prime}_{t+1}$ and $W_{t+1}$:
-$\begin{split}W^{\prime}_{t+1}-W_{t+1}&=-\gamma\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}U_{p,t}\Sigma_{p,t+1}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{train}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\\ &\quad-\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}V^{\top}_{p,t}\\ &\quad+\eta\gamma\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{train}\Sigma_{p,t+1}\\ &\quad+\gamma^{2}\nabla_{U_{p,t}}\mathcal{L}_{reg}\nabla_{V^{\top}_{p,t}}\mathcal{L}_{reg}\Sigma_{p,t+1}\end{split}$ (21)
-Here, we could find that each term has the $\Sigma_{p,t+1}$.
-This shows that by optimizing respected to regularizer loss at step $t+1$ and optimizing $\Sigma_{p,t+1}$ respected to train loss on step $t+2$, we could efficiently increase the orthonormality of $U_{p}$ and $V^{\top}_{p}$ while have negligible side-effect on convergence.
-In short, adding regularizer could transfer scaling information from $U_{p}$ and $V^{\top}_{p}$, which come from optimizing for $\mathcal{L}_{train}$, into $\Sigma_{p}$ allows the matrix scaling information updates more holistically, ultimately presenting a better generalization.
-In Figure 3, we illustrate the effect of $\gamma$ on orthonormality, training loss, and gradient norm.
-The figures present almost no difference in loss and gradient norm in different $\gamma$ values, which indicates the strong robustness of training of SORSA.
+In this section, we provides a gradient analysis for the SORSA with the orthonormal regularizer, focusing on proving the regularizer’s convexity and find the upper bound of the hyperparameter $\gamma$.
+We also demonstrates how the regularizer stabilizes the training and facilitates its generalization.
+To prove SORSA with regularizer is optimizable, we need to prove the convexity of the regularizer, and then use Lipschitz continuity to find the upper bound of $\gamma$.
+In order to analyze the regularizer according to Eq.
+(9), we begin by noticing that the Frobenius norm $\|\cdot\|_{F}$ is a convex function.
+We then examine the terms $U^{\top}_{p}U_{p}-I$ and $V_{p}V^{\top}_{p}-I$, which are quadratic forms in $U_{p}$ and $V_{p}$ , respectively.
+To assess their convexity, we compute the Hessian matrix of each quadratic term with respect to $U_{p}$ and $V_{p}$.
+For the term $U^{\top}_{p}U_{p}-I$ , we first express it as a function of $U_{p}$.
+The Hessian of a quadratic form $U^{\top}_{p}U_{p}$ is given by $2I$, where $I$ is the identity matrix of appropriate dimensions.
+This is because the second-order derivative of $U^{\top}_{p}U_{p}$ with respect to $U_{p}$ is constant and positive semidefinite, ensuring convexity.
+The second term could use the exact same process, which is also convex.
+Therefore, the regularizer is convex.
+Then, we analyze the Lipschitz continuity of the gradients to establish an appropriate bound for $\gamma$. A function $f$ is Lipschitz continuous with constant $L$ if:
+$\|\nabla f(W_{1})-\nabla f(W_{2})\|\leq L\|W_{1}-W_{2}\|$ (12)
+According to Eq. (11), the gradient of combined loss function $\mathcal{L}(W_{p})$ is:
+$\nabla_{W_{p,t}}\mathcal{L}=\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (13)
+Assuming that $\nabla_{W_{p}}\mathcal{L}_{train}$ and $\nabla_{W_{p}}\mathcal{L}_{reg}$ are Lipschitz continuous with constants $L_{train}$ and $L_{reg}$, respectively, the Lipschitz constant for the combined loss function can be written as:
+$L=L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}$ (14)
+When the learning rate is at maximum point, $\eta_{t}=\eta_{d}$, to ensure convergence of the algorithm, the learning rate $\eta_{d}$ must satisfy the following condition:
+$\eta_{d}<\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (15)
+Which could be written as:
+$\gamma<\frac{2-\eta_{d}L_{train}}{L_{reg}}$ (16)
+We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$.
+Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$.
+Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
+$\gamma<\frac{2}{k\eta_{d}}-\frac{L_{train}}{k}$ (17)
+This inequality provides an upper bound on $\gamma$ to ensure that the regularizer does not prevent convergence when $\eta_{t}$ is at its maximum value $\eta_{d}$.
+The inequality shows that $\gamma$ should have inversely proportional relationship to $\eta_{d}$ while selecting the hyperparameter in order to make sure the SORSA adapters could converge.
 Figure 3: Orthonormality, loss and gradient norm variation of SORSA using different $\gamma$ on MetaMathQA training of Mistral 7B v0.1.
+In Figure 3, we could see that below the bound, $\gamma$ have almost no effect on loss curve, reflecting the robustness of SORSA.
+The reinforcement of orthonormality of $U_{p}$ and $V^{\top}_{p}$ concentrates the scaling information into $\Sigma_{p}$, leads to a more holistic parameter updating process.
+The regularizer ensures that $U_{p}$ and $V^{\top}_{p}$ act as pure rotation matrices, while all magnitude updates are captured within $\Sigma_{p}$ by decoupling singular values from singular vectors,
 Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$.
-By ensuring the scaling information only be updated into $\Sigma_{p}$, we could also ensure that the parameter distribution of $W_{p}$ is evenly distributed.
+This separation stabilizes learning, as the singular vectors remain orthonormal, preventing distortions and over-fitting.
+Furthermore, concentrating the updates within $\Sigma_{p}$ allows for a more efficient gradient flow, facilitating smoother updates and ensuring that scaling changes reflect true data variation.
+Consequently, this structured update mechanism fosters better generalization and balanced learning throughout the training process.
 Introducing the orthonormal regularizer into the training process promotes a more stable update trajectory and a more evenly distributed weight by actively countering deviations from orthonormality.
@@
 The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT.
-On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both tasks compared to the other three methods.
+On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both evaluation tasks compared to the other three methods.
 In this section, due to the limitation of computing resources, we train and benchmark a small number of tasks.
@@
 In this paper, we introduced SORSA, a novel parameter-efficient fine-tuning (PEFT) method designed to enhance the adaptation of large language models (LLMs) for downstream tasks.
-SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principle and residual components, only training the principle singular values and vectors while freezing the residuals.
+SORSA utilizes singular value decomposition (SVD) to split pre-trained weights into principal and residual components, only training the principal singular values and vectors while freezing the residuals.
 We implemented an orthonormal regularizer to maintain the orthonormality of singular vectors during training, ensuring efficient parameter updates and preserving the integrity of singular values.
@@
 We also show the significance of the orthonormal regularizer through analysis.
-We also introduce the gradient analysis of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
+We also analyze the regularizer of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
 SORSA retains the advantages of LoRA and variants, including low training VRAM requirements, no inference latency, and versatility across different neural network architectures.
@@
 These experiments could expand SORSA’s applicability to a broader range of deep-learning tasks.
-  • Quantized SORSA (QSORSA): Investigate the application of quantization techniques to SORSA introduced in QLoRA (Dettmers et al., 2024).
+  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024).
 This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources.
@@
 For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training.
-SORSA layers are applied on all linear matrices in every layer.
+SORSA adapters are applied on all linear matrices in every layer.
 We only calculate the loss on the response part.

### v3

@@
-SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models
+SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation
 Abstract
@@
 These parts are initialized by performing SVD on pre-trained weights.
-Moreover, we implement and analyze an orthonormal regularizer, which could effectively transfer the scaling information into $\Sigma_{p}$ and ultimately allows the training process to be more efficient.
+Moreover, we implement and analyze an orthonormal regularizer, which we prove could decrease the condition number of $W_{p}$ and allows the optimization to be more efficient.
 SORSA adapters could be merged during inference, thus eliminating any inference latency.
@@
 The code is available at https://github.com/Gunale0926/SORSA.
-Deep Learning, LLM, PEFT, LoRA, SVD
 Figure 1: Architecture of a SORSA adapter.
@@
 Furthermore, we analyze the pattern of variation of singular values and vectors during parameter updating and discuss the different patterns of partial fine-tuning (FT), LoRA, SORSA without regularizer, and SORSA with regularizer concerning singular values and vectors’ updating.
+We also provide a comprehensive gradient analysis that provides a mathematical foundation for SORSA.
+This analysis demonstrates several crucial properties of our method, including the convexity of the regularizer, Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$.
+Moreover, we prove that SORSA improves the condition number of the optimization problem compared to unregularized approaches.
 In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
@@
 This approach significantly reduces the memory and computational requirements during training, especially VRAM.
-Adapter-based PEFT is the first type of PEFT, which was initially designed by (Houlsby et al., 2019).
+Adapter-based PEFT is the first type of PEFT, which was initially designed by Houlsby et al.
+(2019).
 It introduces additional trainable non-linear blocks into the frozen pre-trained model, which could effectively tune the pre-trained model with a limited amount of trainable parameters.
-Its variants, e.g., (Lin et al., 2020), reduce the number of adapter layers per block, and (He et al., 2022) focus on adding adapter modules parallel to existing layers.
+Its variants, e.g., Lin et al.
+(2020), reduce the number of adapter layers per block, and He et al.
+(2022) focus on adding adapter modules parallel to existing layers.
 However, all adapter-based PEFT methods introduce inference latency due to their non-mergeable attribute.
-Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in (Lester et al., 2021).
-There are several variants of this work, including (Liu et al., 2022; Razdaibiedina et al., 2023).
+Prompt-based PEFT is also a well-known type of PEFT, which was first proposed in Lester et al.
+(2021).
+There are several variants of this work, including Liu et al.
+(2022); Razdaibiedina et al.
+(2023).
 However, they have some inevitable shortcomings, such as potential performance limitations compared to full parameter fine-tuned models, additional inference latency due to expanding the length of the total input to the model, and the complexity of designing effective initialization.
@@
 This property allows SORSA adapters to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
-4 Our Method
-4.1 Singular Value and Orthonormal Regularized Singular Vector Adaptation
+4 Singular Value and Orthonormal Regularized Singular Vector Adaptation
 According to the definition of SVD in Section 3.1, given a rank $r$ where $r\ll k$, we could perform the low-rank approximation by selecting the first $r$ items on the diagonal of $\Sigma$, which is the first $r$ most significant singular values, and also select the first $r$ columns of $U$ and first $r$ rows of $V^{\top}$, which correspond to the selected singular values.
@@
 Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principal weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
-$W_{0}=W_{p}+W_{r}$ (5)
 $W_{p}=U_{[:,:r]}\Sigma_{[:r,:r]}V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
@@
 Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices.
-SORSA is defined by Eq (8), which is initially equivalent to the pre-trained weight $W_{0}$.
+SORSA is defined by Eq equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
 During training, $W_{r}$ remains frozen, and only $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ are updated.
@@
 $\text{SORSA}(x):=W_{r}x+W_{p}x$ (7)
-We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 4.2.
-$\mathcal{L}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V_{p}V^{\top}_{p}-I\|_{F}$ (9)
+We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 5.
+${\mathcal{L}}_{reg}=\|U^{\top}_{p}U_{p}-I\|_{F}+\|V^{\top}_{p}V_{p}-I\|_{F}$ (8)
 Where ${\mathcal{L}}_{reg}$ is the orthonormal regularizer loss, the $U_{p}$ and $V^{\top}_{p}$ are each orthonormal vectors in columns and rows, respectively, after initialization due to SVD’s property.
@@
 $W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{train}-\gamma_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ (9)
-At training step $t$, $\nabla_{W_{t}}\mathcal{L}_{train}$ denotes the gradient of $W_{t}$ respect to $\mathcal{L}_{train}$, and $\nabla_{W_{t}}\mathcal{L}_{reg}$ denotes the gradient of $W_{t}$ respect to the orthonormal regularizer loss $\mathcal{L}_{reg}$.
-$\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss, respectively.
-We update the SORSA as Eq (11) for implementation simplicity.
+At training step $t$, $\nabla_{W_{p,t}}{\mathcal{L}}_{train}$ denotes the gradient of ${\mathcal{L}}_{train}$ respect to $W_{p,t}$, and $\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ denotes the gradient of ${\mathcal{L}}_{reg}$ respect to the orthonormal regularizer loss $W_{p,t}$.
+$\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss at step $t$, respectively.
+We update the SORSA as Eq equation 10 for implementation simplicity.
 $W_{p,t+1}=$ $W_{p,t}-\eta_{t}\left(\nabla_{W_{p,t}}{\mathcal{L}}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}\right)$ (10)
-$\eta_{d}$ is the learning rate which is given to the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
-4.2 Gradient Analysis
-In this section, we provides a gradient analysis for the SORSA with the orthonormal regularizer, focusing on proving the regularizer’s convexity and find the upper bound of the hyperparameter $\gamma$.
-We also demonstrates how the regularizer stabilizes the training and facilitates its generalization.
-To prove SORSA with regularizer is optimizable, we need to prove the convexity of the regularizer, and then use Lipschitz continuity to find the upper bound of $\gamma$.
-In order to analyze the regularizer according to Eq.
-(9), we begin by noticing that the Frobenius norm $\|\cdot\|_{F}$ is a convex function.
-We then examine the terms $U^{\top}_{p}U_{p}-I$ and $V_{p}V^{\top}_{p}-I$, which are quadratic forms in $U_{p}$ and $V_{p}$ , respectively.
-To assess their convexity, we compute the Hessian matrix of each quadratic term with respect to $U_{p}$ and $V_{p}$.
-For the term $U^{\top}_{p}U_{p}-I$ , we first express it as a function of $U_{p}$.
-The Hessian of a quadratic form $U^{\top}_{p}U_{p}$ is given by $2I$, where $I$ is the identity matrix of appropriate dimensions.
-This is because the second-order derivative of $U^{\top}_{p}U_{p}$ with respect to $U_{p}$ is constant and positive semidefinite, ensuring convexity.
-The second term could use the exact same process, which is also convex.
-Therefore, the regularizer is convex.
-Then, we analyze the Lipschitz continuity of the gradients to establish an appropriate bound for $\gamma$. A function $f$ is Lipschitz continuous with constant $L$ if:
-$\|\nabla f(W_{1})-\nabla f(W_{2})\|\leq L\|W_{1}-W_{2}\|$ (12)
-According to Eq. (11), the gradient of combined loss function $\mathcal{L}(W_{p})$ is:
-$\nabla_{W_{p,t}}\mathcal{L}=\nabla_{W_{p,t}}\mathcal{L}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}\mathcal{L}_{reg}$ (13)
-Assuming that $\nabla_{W_{p}}\mathcal{L}_{train}$ and $\nabla_{W_{p}}\mathcal{L}_{reg}$ are Lipschitz continuous with constants $L_{train}$ and $L_{reg}$, respectively, the Lipschitz constant for the combined loss function can be written as:
-$L=L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}$ (14)
-When the learning rate is at maximum point, $\eta_{t}=\eta_{d}$, to ensure convergence of the algorithm, the learning rate $\eta_{d}$ must satisfy the following condition:
-$\eta_{d}<\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (15)
-Which could be written as:
-$\gamma<\frac{2-\eta_{d}L_{train}}{L_{reg}}$ (16)
-We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$.
-Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$.
-Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
-$\gamma<\frac{2}{k\eta_{d}}-\frac{L_{train}}{k}$ (17)
-This inequality provides an upper bound on $\gamma$ to ensure that the regularizer does not prevent convergence when $\eta_{t}$ is at its maximum value $\eta_{d}$.
-The inequality shows that $\gamma$ should have inversely proportional relationship to $\eta_{d}$ while selecting the hyperparameter in order to make sure the SORSA adapters could converge.
-Figure 3: Orthonormality, loss and gradient norm variation of SORSA using different $\gamma$ on MetaMathQA training of Mistral 7B v0.1.
-In Figure 3, we could see that below the bound, $\gamma$ have almost no effect on loss curve, reflecting the robustness of SORSA.
-The reinforcement of orthonormality of $U_{p}$ and $V^{\top}_{p}$ concentrates the scaling information into $\Sigma_{p}$, leads to a more holistic parameter updating process.
-The regularizer ensures that $U_{p}$ and $V^{\top}_{p}$ act as pure rotation matrices, while all magnitude updates are captured within $\Sigma_{p}$ by decoupling singular values from singular vectors,
-Moreover, training along with an Adam-like (Kingma & Ba, 2015) optimizer could accelerate this process because its momentum on parameters of $U_{t}$ will become much smaller than momentum on parameters of $\Sigma_{t}$, which means the optimizer will halt the parameter updating towards the direction point to orthonormality increasing in $U_{p}$ and $V^{\top}_{p}$, and more likely to update the scaling information to $\Sigma_{p}$.
-This separation stabilizes learning, as the singular vectors remain orthonormal, preventing distortions and over-fitting.
-Furthermore, concentrating the updates within $\Sigma_{p}$ allows for a more efficient gradient flow, facilitating smoother updates and ensuring that scaling changes reflect true data variation.
-Consequently, this structured update mechanism fosters better generalization and balanced learning throughout the training process.
-Introducing the orthonormal regularizer into the training process promotes a more stable update trajectory and a more evenly distributed weight by actively countering deviations from orthonormality.
-This regulation stabilizes the updates and enhances the model’s ability to maintain crucial properties of the singular vectors $U_{p}$ and $V^{\top}_{p}$, thereby contributing to improved generalization and effectiveness in learning.
-This analysis underlines the utility of orthonormal regularization in complex model training scenarios, particularly in fine-tuning pre-trained models where preserving learned features is crucial.
+$\eta_{d}$ is the maximum learning rate which is given from the scheduler. This implementation allows us to use only one optimizer and scheduler to deal with two different learning rates separately.
+5 Gradient Analysis
+In this section, we present a comprehensive mathematical analysis of the SORSA method which mainly focus on the effect of orthonormal regularization.
+Our investigation elucidates the key optimization properties of SORSA, providing a theoretical foundation for its advantages.
+We explore four critical aspects: the convexity of the regularizer, the Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$, and the impact on the condition number of the optimization problem.
+The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix B.
+Our analysis reveals fundamental theoretical properties of SORSA, establishing its mathematical soundness and demonstrating its optimization advantages.
+We prove two key theorems that form the cornerstone of our theoretical framework:
+Theorem 5.1.
+The regularizer ${\mathcal{L}}_{reg}$ is convex.
+Theorem 5.2.
+The gradient of the regularizer ${\mathcal{L}}_{reg}$ is Lipschitz continuous.
+These properties collectively ensure a well-behaved optimization landscape, facilitating stable and efficient training of SORSA adapters.
+Building upon these foundational results, we further analyze the bounds of the hyperparameter $\gamma$, a critical factor in the performance of SORSA:
+Theorem 5.3.
+For convergence of gradient descent, the learning rate $\eta_{d}$ and regularization parameter $\gamma$ must satisfy:
+$\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
+This theorem provides crucial guidance for practitioners, offering a clear criterion for selecting appropriate values of $\gamma$ to ensure convergence of the gradient descent process.
+To demonstrate the superior optimization properties of SORSA, we present a novel analysis of its condition number, a critical factor in determining convergence speed and stability.
+Our theoretical investigation reveals a significant improvement in the condition number compared to unregularized approaches, providing a mathematical foundation for SORSA’s enhanced performance.
+We begin this analysis by establishing a key lemma that bounds the effect of the orthonormal regularizer on the singular values of the weight matrix:
+Lemma 5.4.
+Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\Sigma_{p}^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\Sigma_{p}^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer.
+For each singular value $\sigma_{i}$, the following bound holds:
+$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (12)
+where $\epsilon$ is a small positive constant.
+Lemma 5.4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
+Theorem 5.5.
+The orthonormal regularizer in SORSA can improve the condition number of the optimization problem over the course of training under certain conditions. Specifically:
+At initialization ($t=0$):
+$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=1$ (13)
+When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$.
+When $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
+This theorem quantifies the improvement in the condition number achieved by SORSA, offering an explanation for its fast convergence.
+The proof leverages the effects of the orthonormal regularization to establish a tight bound on the condition number ratio.
+This theorem could also shows that training with the regularizer, the distribution of $W_{p}$ will be more evenly distributed due to a smaller ratio between $\sigma_{\max}(W_{p})$ and $\sigma_{\min}(W_{p})$, which means a better training stability.
+In conclusion, these theorems provide a mathematical foundation for the SORSA method.
+These theoretical guarantees validate the empirical success of SORSA but also provide valuable insights for future developments in PEFT methods.
 6 Empirical Experiments
@@
 We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023).
-We trained the models with SORSA applied and conducted GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) tests.
+We trained the models with SORSA applied using MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
+We also trained the model on CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021).
 The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024).
+All reported values are accuracy.
 See Appendix A.2 for more details and hyperparameters of the training.
-We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from the PiSSA paper (Meng et al., 2024).
-For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80G) GPU.
+We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from Meng et al.
+(2024).
+For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80GB) GPU.
 See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
-See Figure 5 for the variation in orthonormality of SORSA.
-Model Method Trainable Parameters MATH GSM-8K
-Llama 2 7B Full FT 6738M 7.22 49.05
-LoRA 320M 5.50 42.30
-PiSSA 320M 7.44 53.07
-SORSA 320M 10.36 56.03
-Mistral 7B Full FT 7242M 18.60 67.02
-LoRA 168M 19.68 67.70
-PiSSA 168M 21.54 72.86
-SORSA 168M 21.86 73.09
-Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from the paper of PiSSA (Meng et al., 2024).
+Model Method Trainable Parameters MATH GSM-8K HumanEval
+Llama 2 7B Full FT 6738M 7.22 49.05 21.34
+LoRA 320M 5.50 42.30 18.29
+PiSSA 320M 7.44 53.07 21.95
+SORSA 320M 10.36 56.03 24.39
+Mistral 7B Full FT 7242M 18.60 67.02 45.12
+LoRA 168M 19.68 67.70 43.90
+PiSSA 168M 21.54 72.86 46.95
+SORSA 168M 21.86 73.09 47.56
+Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from Meng et al. (2024).
 Refer to caption Refer to caption
-Figure 4: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B.
-Figure 5: Variation of orthonormality on MetaMathQA training of Llama 2 7B and Mistral 7B v0.1.
+Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B. LoRA and PiSSA curves are from Meng et al. (2024).
 The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT.
@@
 We also show the significance of the orthonormal regularizer through analysis.
-We also analyze the regularizer of SORSA in the paper, which demonstrates the process of scaling information to be transferred into $\Sigma_{p}$ during training, which makes the parameter updating more effective because it directly changes the global characteristics of the matrix.
+Our gradient analysis provides a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape.
+This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
 SORSA retains the advantages of LoRA and variants, including low training VRAM requirements, no inference latency, and versatility across different neural network architectures.
@@
 While SORSA demonstrates substantial improvements over existing PEFT methods, there are several avenues for future work to enhance its capabilities further and extend its applicability:
-  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different transformer-based LLMs.
-This includes replicating other tests from studies like PiSSA (Meng et al., 2024) and performing evaluations such as HumanEval (Chen et al., 2021) and MT-Bench (Zheng et al., 2023), in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
-  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2021a) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
-  • Application to Other Fields: Test SORSA on different fields beyond NLP.
-Such as conducting experiments on computer vision (CV) models like DDPM (Ho et al., 2020), DDIM (Song et al., 2020), and ViT (Dosovitskiy et al., 2021).
-These experiments could expand SORSA’s applicability to a broader range of deep-learning tasks.
+  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different models, in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
+  • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2020) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
   • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024).
@@
 The PEFT method, while efficient, raises critical ethical concerns regarding the security of built-in safety measures in AI models.
-As demonstrated in (Lermen & Rogers-Smith, 2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content.
+As demonstrated in Lermen & Rogers-Smith (2024), subversive fine-tuning techniques can bypass safety training intended to prevent the generation of harmful content.
 The ease and affordability of such methods underscore the vulnerability of safety protocols.
@@
 We would like to extend our gratitude to Prof.
-Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing guidance to the research.
-Appendix A Training Details
+Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing the endorsement to the research.
+Appendix A Experiments Details
 A.1 Analysis
-For the singular values and vectors analysis in section 3, we apply partial fine-tuning, LoRA SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, using the same training data and optimizer as A.2.
-In the analysis, LoRA and SORSA are only applied to q_proj and v_proj matrices, respectively.
-FT is only trainable on q_proj and v_proj.
-We only calculate the loss on the response part.
-The models are trained with BF16 mix precision.
+For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset.
+In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively.
+For Partial FT, we set model’s q_proj and v_proj matrices to trainable.
+We only calculated the loss on the response part.
+The models are trained with TF32 + BF16 mix precision.
 See Table 2 for hyperparameters.
+We also found we should only perform SVD for analysis using CPU, in order to get the precise analysis data.
 Training
@@
 A.2 NLG Experiments
-For our NLG tasks, we adapt Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by LoRA (Hu et al., 2021), PiSSA (Meng et al., 2024) and SORSA, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset, and test by GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021b) dataset, We use AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training.
-SORSA adapters are applied on all linear matrices in every layer.
-We only calculate the loss on the response part.
-The models are trained with BF16 mix precision.
+For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by SORSA.
+For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset.
+For HumanEval (Chen et al., 2021) evaluation, we use the first 100K data in CodeFeedback Filtered Instruction(Zheng et al., 2024) dataset.
+We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training.
+SORSA adapters were applied on all linear matrices in every layer.
+We only calculated the loss on the response part.
+The models are trained with TF32 + BF16 mix precision.
 See Table 3 for hyperparameters.
@@
 MATH Max Length 2048
+HumanEval Max Length 2048
 Table 3: Hyperparameters of training on models for GSM-8K and MATH
+1@@ Instruction
+2Here is the given code to do completion:
+3‘‘‘python
+4{prompt}
+5‘‘‘
+6
+7Please continue to complete the function with python programming language. You are not allowed to modify the given code and do the completion only.
+8
+9Please return all completed codes in one code block.
+10This code block should be in the following format:
+11’’’python
+12# Your codes here
+13’’’
+14
+15@@ Response
+Listing 1: Prompt used for HumanEval evaluation.
+Appendix B Proofs
+B.1 Proof of Theorem 5.1
+Proof.
+We prove this in two steps:
+First, we show that $f(U_{p})=\|U^{\top}_{p}U_{p}-I\|_{F}$ is convex. Then, we prove that $g(V_{p})=\|V_{p}V^{\top}_{p}-I\|_{F}$ is convex.
+Since the sum of convex functions is convex, this will establish the convexity of ${\mathcal{L}}_{reg}$.
+Let $U_{p},W\in\mathbb{R}^{m\times r}$. The Hessian of $f$ at $U_{p}$ in the direction $W$ is given by:
+$\begin{split}\nabla^{2}f(U_{p})[W,W]&=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(f(U_{p}+\epsilon W)-2f(U_{p})+f(U_{p}-\epsilon W)\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|(U_{p}+\epsilon W)^{\top}(U_{p}+\epsilon W)-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|(U_{p}-\epsilon W)^{\top}(U_{p}-\epsilon W)-I\|_{F}\Big)\\ &=\lim_{\epsilon\to 0}\frac{1}{\epsilon^{2}}\Big(\|U^{\top}_{p}U_{p}+\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\\ &\quad-2\|U^{\top}_{p}U_{p}-I\|_{F}+\|U^{\top}_{p}U_{p}-\epsilon(U^{\top}_{p}W+W^{\top}U_{p})+\epsilon^{2}W^{\top}W-I\|_{F}\Big)\\ &=2\|W^{\top}W\|_{F}\end{split}$ (14)
+Since $\|W^{\top}W\|_{F}\geq 0$ for all $W$, we have $\nabla^{2}f(U_{p})[W,W]\geq 0$, which proves that $f$ is convex.
+The proof for $g(V_{p})$ follows the same steps as for $f(U_{p})$, leading to the same conclusion.
+Therefore, both $f(U_{p})$ and $g(V_{p})$ are convex, and consequently, ${\mathcal{L}}_{reg}$ is convex. ∎
+B.2 Proof of Theorem 5.2
+Proof.
+To prove Lipschitz continuity, we need to show that there exists a constant $L>0$ such that for any two pairs of matrices $(U_{1},V_{1})$ and $(U_{2},V_{2})$:
+$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (15)
+Let’s break this down into two parts:
+First, consider $|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|$:
+$\begin{split}|\|U_{1}^{\top}U_{1}-I\|_{F}-\|U_{2}^{\top}U_{2}-I\|_{F}|&\leq\|(U_{1}^{\top}U_{1}-I)-(U_{2}^{\top}U_{2}-I)\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{2}^{\top}U_{2}\|_{F}\\ &=\|U_{1}^{\top}U_{1}-U_{1}^{\top}U_{2}+U_{1}^{\top}U_{2}-U_{2}^{\top}U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}(U_{1}-U_{2})\|_{F}+\|(U_{1}^{\top}-U_{2}^{\top})U_{2}\|_{F}\\ &\leq\|U_{1}^{\top}\|_{F}\|U_{1}-U_{2}\|_{F}+\|U_{1}-U_{2}\|_{F}\|U_{2}\|_{F}\\ &\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\end{split}$ (16)
+Here, we’ve used the triangle inequality and the sub-multiplicative property of the Frobenius norm.
+Similarly for $V$:
+$|\|V_{1}^{\top}V_{1}-I\|_{F}-\|V_{2}^{\top}V_{2}-I\|_{F}|\leq(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}$ (17)
+Combining these results:
+$\begin{split}|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|&\leq(\|U_{1}\|_{F}+\|U_{2}\|_{F})\|U_{1}-U_{2}\|_{F}\\ &\quad+(\|V_{1}\|_{F}+\|V_{2}\|_{F})\|V_{1}-V_{2}\|_{F}\\ &\leq\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})\\ &\quad\cdot(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})\end{split}$ (18)
+Let $L_{reg}=\max(\|U_{1}\|_{F}+\|U_{2}\|_{F},\|V_{1}\|_{F}+\|V_{2}\|_{F})$.
+This $L$ is finite because the Frobenius norms of $U$ and $V$ are bounded (they represent orthonormal matrices in the ideal case).
+Therefore, we have shown that:
+$|{\mathcal{L}}_{reg}(U_{1},V_{1})-{\mathcal{L}}_{reg}(U_{2},V_{2})|\leq L_{reg}(\|U_{1}-U_{2}\|_{F}+\|V_{1}-V_{2}\|_{F})$ (19)
+This proves that ${\mathcal{L}}_{reg}$ is Lipschitz continuous with Lipschitz constant $L_{reg}$. ∎
+B.3 Proof of Theorem 5.3
+Proof.
+For gradient descent to converge while the $\eta_{t}=\eta_{d}$, we need:
+$\eta_{d}<\frac{2}{L}=\frac{2}{L_{train}+\frac{\gamma}{\eta_{d}}L_{reg}}$ (20)
+Rearranging this inequality:
+$\begin{split}\eta_{d}(L_{train}+\frac{\gamma}{\eta_{d}}L_{reg})&<2\\ \eta_{d}L_{train}+\gamma L_{reg}&<2\\ \gamma L_{reg}&<2-\eta_{d}L_{train}\\ \gamma&<\frac{2-\eta_{d}L_{train}}{L_{reg}}\end{split}$ (21)
+We can assume that the regularizer’s gradients scale with $\eta_{d}$ , meaning that larger updating step (due to a larger $\eta_{d}$) will lead to greater deviations from orthonormality, which increases $L_{reg}$.
+Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$.
+Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
+$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (22)
+This establishes the upper bound for $\gamma$. ∎
+B.4 Proof of Lemma 5.4
+Proof.
+First, let’s consider the effect of the orthonormal regularizer. The regularizer aims to make $U_{p}^{\top}U_{p}\approx I$ and $V_{p}V_{p}^{\top}\approx I$. We can quantify this approximation as:
+$\nabla_{W_{p}}{\mathcal{L}}_{reg}\leq\epsilon_{\nabla}$ (23)
+where $\epsilon_{\nabla}>0$ is a small constant.
+Now, let’s consider the difference between $W_{p}^{\text{unreg}}$ and $W_{p}^{\text{reg}}$:
+$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}$ (24)
+Now, we can use Weyl’s inequality, which states that for matrices A and B:
+$|\sigma_{i}(A+B)-\sigma_{i}(A)|\leq\|B\|_{2}\leq\|B\|_{F}$ (25)
+Applying this to our case, with $A=W_{p}^{\text{unreg}}$ and $B=W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}$:
+$|\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}|\leq\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}\leq\gamma\epsilon_{\nabla}$ (26)
+Let $\epsilon=\gamma\epsilon_{\nabla}$. Then we have:
+$-\epsilon\leq\sigma_{i}^{\text{reg}}-\sigma_{i}^{\text{unreg}}\leq\epsilon$ (27)
+Rearranging this inequality gives us our desired bound:
+$(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (28)
+B.5 Proof of Theorem 5.5
+Proof.
+Let $W_{p}=U_{p}\Sigma_{p}V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
+$\kappa(W_{p})=\frac{\sigma_{\max}(W_{p})}{\sigma_{\min}(W_{p})}$ (29)
+where $\sigma_{\max}$ and $\sigma_{\min}$ are the maximum and minimum singular values of $W_{p}$.
+At initialization ($t=0$): Due to SVD initialization, $U_{p}$ and $V_{p}^{\top}$ are perfectly orthonormal, so:
+$\kappa(U_{p}^{\text{unreg}})=\kappa((V_{p}^{\text{unreg}})^{\top})=\kappa(U_{p}^{\text{reg}})=\kappa((V_{p}^{\text{reg}})^{\top})=1$ (30)
+And $\epsilon_{0}=0$, $\delta_{1,0}=\delta_{2,0}=0$. Therefore:
+$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\Sigma_{p}^{\text{reg}})}{\kappa(\Sigma_{p}^{\text{unreg}})}=1$ (31)
+During training ($t>0$): As training progresses, $U_{p}$ and $V_{p}^{\top}$ deviate from orthonormality in the unregularized case. We quantify this deviation:
+$\|U_{p}^{\top}U_{p}-I\|_{F}\leq\epsilon_{1,t}$ (32)
+$\|V_{p}^{\top}V_{p}-I\|_{F}\leq\epsilon_{2,t}$ (33)
+where $\epsilon_{1,t},\epsilon_{2,t}>0$ increase over time t.
+For the regularized matrices, we can bound their condition numbers:
+$\kappa(U_{p}^{\text{reg}})\leq 1+\delta_{1,t}$ (34)
+$\kappa(V_{p}^{\text{reg}})\leq 1+\delta_{2,t}$ (35)
+where $\delta_{1,t},\delta_{2,t}$ are small positive numbers that remain bounded due to the regularization.
+From the Lemma 5.4, we arrive at:
+$\begin{split}\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1-\epsilon_{t}}{1+\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\leq\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}\leq\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\end{split}$ (36)
+As training continues, in the unregularized case, $\kappa(U_{p,\text{unreg}})$ and $\kappa(V_{p,\text{unreg}}^{\top})$ tend to increase as $U_{p}$ and $V_{p}^{\top}$ deviate further from orthonormality.
+On the other hand, $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will approach to 1 because of the reinforcement in orthonormality.
+Therefore, $\exists c>0$, while $t>c$:
+$\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})>(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ (37)
+will hold true.
+Therefore, while $t>c$, we have:
+$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}<1$ (38)
+indicating an improvement in the condition number. ∎

### v4

@@
-SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation
+SORSA: Singular Values and Orthonormal Regularized Singular Vectors Adaptation of Large Language Models
 Abstract
@@
 We introduce a method to analyze the variation of the parameters by performing singular value decomposition (SVD) and discuss and analyze SORSA’s superiority in minimizing the alteration in the SVD aspect.
-Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
+Each SORSA adapter consists of two main parts: trainable principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and frozen residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$.
 These parts are initialized by performing SVD on pre-trained weights.
@@
 After all, SORSA shows a faster convergence than PiSSA and LoRA in our experiments.
-On the MATH benchmark, Llama 2 7B adapted using SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%).
-On the GSM-8K benchmark, SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%).
+On the GSM-8K benchmark, Llama 2 7B adapted using SORSA achieved 56.03% accuracy, surpassing LoRA (42.30%), Full FT (49.05%), and PiSSA (53.07%).
+On the MATH benchmark, SORSA achieved 10.36% accuracy, outperforming LoRA (5.50%), Full FT (7.22%), and PiSSA (7.44%).
 We conclude that SORSA offers a new perspective on parameter-efficient fine-tuning, demonstrating remarkable performance.
@@
 Figure 1: Architecture of a SORSA adapter.
-We only train parts rendered in orange ($U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$).
+We only train parts rendered in orange ($U_{p}$, $\text{diag}(S_{p})$ and $V^{\top}_{p}$), and freeze parts rendered in blue ($U_{r}$, $\text{diag}(S_{r})$ and $V^{\top}_{r}$).
 1 Introduction
-Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks.
+Pre-trained large language models (LLMs) show remarkable generalization abilities, allowing them to perform various kinds of natural language processing (NLP) tasks (Peng et al., 2024; Touvron et al., 2023; Dubey et al., 2024; Radford et al., 2019; OpenAI, 2023).
 For specific downstream tasks, full parameter fine-tuning, which continues training all parameters of LLMs on downstream data, is widely used.
@@
 For example, the estimated VRAM requirement for fully fine-tuning Llama 2 7B using Float32 could approach approximately 100 GB, making it unlikely to fully fine-tune the model on a single GPU with current technology.
-Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB, thus rendering it unfeasible on a single GPU with current technology.
+Additionally, the VRAM requirement for fully fine-tuning Llama 2 70B using Float32 exceeds 1 TB (Touvron et al., 2023; Anthony et al., 2023), thus rendering it unfeasible on a single GPU with current technology.
 To address these challenges, several parameter-efficient fine-tuning (PEFT) methods (Houlsby et al., 2019; Lester et al., 2021; Hu et al., 2021) have been proposed.
 These methods enable the training of only a limited number of parameters, which significantly reduces VRAM requirements while achieving comparable or even superior performance to full fine-tuning.
-For instance, tuning Llama 2 7B in Float32 by LoRA with a rank of 128 only takes approximately 60GB VRAM, which allows training in 1 $\times$ NVIDIA A100-80GB or 3 $\times$ NVIDIA RTX 4090-24GB.
+For instance, tuning Llama 2 7B in Float32 by LoRA (Hu et al., 2021) with a rank of 128 only takes approximately 60GB VRAM, which allows training on 1 $\times$ NVIDIA A100 (80GB), or even 3 $\times$ NVIDIA RTX 4090 (24GB).
 Among those PEFT methods, LoRA (Hu et al., 2021) and its variants (Zhang et al., 2023; Meng et al., 2024; Liu et al., 2024; Dettmers et al., 2024) had become increasingly popular due to their: 1.
@@
 In this paper, we propose a novel PEFT approach, Singular Values and Orthonormal Regularized Singular Vectors Adaptation, or SORSA.
-A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\Sigma_{p}V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\Sigma_{r}V^{\top}_{r}$.
+A SORSA adapter has two main parts: principal singular weights $W_{p}=U_{p}\text{diag}(S_{p})V^{\top}_{p}$, and residual weights $W_{r}=U_{r}\text{diag}(S_{r})V^{\top}_{r}$.
 These two parts are initialized by performing singular value decomposition (SVD) on pre-trained weight.
@@
 Moreover, we prove that SORSA improves the condition number of the optimization problem compared to unregularized approaches.
-In our experiments, SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning.
+SORSA retains all the benefits of LoRA and its variants while demonstrating remarkable performance compared to PiSSA, LoRA, and full parameter fine-tuning in our experiments.
 2 Related Works
@@
 However, DoRA introduced a calculation of norm in every training step, which makes it much more inefficient in training than LoRA.
-PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=BA$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen.
+OLoRA (Büyükakyüz, 2024) uses QR decomposition to initialize the LoRA adapters $A$ and $B$, which initializes $B$ as an orthogonal matrix.
+They discusses the significance of orthonormality in neural networks’ weight (See Section 5 for more details).
+In their experiments, OLoRA demonstrates faster convergence than LoRA.
+PiSSA (Meng et al., 2024) decomposes pre-trained weight $W_{0}=U\text{diag}(S)V^{\top}$ by Singular Value Decomposition (SVD), and then splits $W_{0}$ into $W_{pri}$ and $W_{res}$: $W_{pri}=AB$ which is trainable, where $A=U_{[:r]}\text{diag}(S^{\frac{1}{2}}_{[:r]})$ and $B=\text{diag}(S^{\frac{1}{2}}_{[:r]})V^{\top}_{[:r]}$; $W_{res}=U_{[:r,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}$ which is frozen.
 PiSSA results a faster convergence speed and better fitting compared to LoRA.
@@
 In other words, the matrix $W$ can be represented as $\Sigma$ on these selected bases, where $\Sigma$ is a diagonal matrix with all non-negative diagonal elements.
-According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\Sigma V^{\top}$.
-Here, $U\in\mathbb{R}^{m\times k}$ are left singular vectors and have orthonormal columns, $V^{\top}\in\mathbb{R}^{k\times n}$ are right singular vectors and have orthonormal rows, and $\Sigma\in\mathbb{R}^{k\times k}$ is a diagonal matrix, where diagonal values are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order.
+According to the meaning of SVD, giving a matrix $W\in\mathbb{R}^{m\times n}$, let $k=\min(m,n)$, we could perform SVD to decompose $W$ by $W=U\text{diag}(S)V^{\top}$.
+Here, $U\in\mathbb{R}^{m\times k}$ is matrix of left singular vectors and have orthonormal columns, $V\in\mathbb{R}^{n\times k}$ is matrix of right singular vectors and have orthonormal columns, and $S\in\mathbb{R}^{k}$ are singular values $\sigma^{1},\sigma^{2}\ldots\sigma^{k}$ arranged in descending order.
+$\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$.
 3.2 Analysis Method
@@
 Our analysis suggests a significant difference in singular values and vectors’ stability and an updating pattern of partial fine-tuning, LoRA, and SORSA.
-The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m,n}$ and tuned weight $W_{t}\in\mathbb{R}^{m,n}$, which $t$ denotes the training step, could be defined as follows:
+The singular value and vector variations between pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$ and tuned weight $W_{t}\in\mathbb{R}^{m\times n}$, which $t$ denotes the training step, could be defined as follows:
 $\Delta\Sigma_{t}=\frac{\Sigma_{i=1}^{k}\left|\sigma_{t}^{i}-\sigma_{0}^{i}\right|}{k}$ (1)
@@
 We test partial fine-tuning, LoRA, and SORSA (with regularizer and without regularizer).
-See Appendix A.1 for training details of the analysis.
+See Section B.1 for training details of the analysis.
 3.3 Analysis Result
-This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect.
-The analysis data is illustrated in Figure 2.
-The four methods obtained similar final loss ($\pm 10\%$).
+This section analyzes the results of different training methods: partial fine-tuning, LoRA, and SORSA based on the data we collect. The analysis data is illustrated in Figure 2.
 Refer to caption Refer to caption
@@
 This property allows SORSA adapters to be trained with higher learning rates without showing noticeable over-fitting compared to other tuning methods.
-4 Singular Value and Orthonormal Regularized Singular Vector Adaptation
+4 SORSA: Singular Value and Orthonormal Regularized Singular Vector Adaptation
 According to the definition of SVD in Section 3.1, given a rank $r$ where $r\ll k$, we could perform the low-rank approximation by selecting the first $r$ items on the diagonal of $\Sigma$, which is the first $r$ most significant singular values, and also select the first $r$ columns of $U$ and first $r$ rows of $V^{\top}$, which correspond to the selected singular values.
@@
 Therefore, for a pre-trained weight $W_{0}\in\mathbb{R}^{m\times n}$, we could split it based on its singular value into principal weight $W_{p}$ and residual weight $W_{r}$, where $W_{p}$ contains the most important part of information of the matrix, and $W_{r}$ contains the least significant part:
-$W_{p}=U_{[:,:r]}\Sigma_{[:r,:r]}V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
-$W_{r}=U_{[:,r:]}\Sigma_{[r:,r:]}V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (6)
-Here, $U$ represents the left singular vectors, $\Sigma$ represents the diagonal matrix of singular values, and $V^{\top}$ represents the right singular vectors.
+$W_{p}=U_{[:,:r]}\text{diag}(S_{[:r]})V^{\top}_{[:r,:]}\in\mathbb{R}^{m\times n}$ (5)
+$W_{r}=U_{[:,r:]}\text{diag}(S_{[r:]})V^{\top}_{[r:,:]}\in\mathbb{R}^{m\times n}$ (6)
+Here, $U$ represents the matrix of left singular vectors, $S$ represents the singular values, $\text{diag}(W)$ denotes a function to form a diagonal matrix from $W$, and $V$ represents the matrix of right singular vectors.
 We use PyTorch (Paszke et al., 2019) syntax to demonstrate matrix selection, where $[:,:r]$ denotes selecting the first $r$ columns of the matrix, and $[r:,:]$ denotes selecting the last $r$ rows of the matrix.
-We rewrite $U_{[:,:r]}$, $\Sigma_{[:r,:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $\Sigma_{[r:,r:]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $\Sigma_{r}$ and $V^{\top}_{r}$ correspondingly.
+We rewrite $U_{[:,:r]}$, $S_{[:r]}$ and $V^{\top}_{[:r,:]}$, which consist $W_{p}$ as $U_{p}$, $S_{p}$ and $V^{\top}_{p}$ for simplicity, and rewrite $U_{[:,r:]}$, $S_{[:r]}$ and $V^{\top}_{[r:,:]}$, which consist $W_{r}$ as $U_{r}$, $S_{r}$ and $V^{\top}_{r}$ correspondingly.
 The initialization of $W_{r}$ in SORSA is the same as PiSSA (Meng et al., 2024).
-Nevertheless, unlike PiSSA which merge $\Sigma_{p}$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\Sigma_{p}^{\frac{1}{2}}$ and $B=\Sigma_{p}^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ in separate matrices.
-SORSA is defined by Eq equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
-During training, $W_{r}$ remains frozen, and only $U_{p}$, $\Sigma_{p}$, and $V^{\top}_{p}$ are updated.
-Since $\Sigma_{p}$ is a diagonal matrix, only the diagonal elements are updated, while the rest of the matrix remains zero.
-$\text{SORSA}(x):=W_{r}x+W_{p}x$ (7)
+Nevertheless, unlike PiSSA which merge $\text{diag}(S_{p})$ with $U_{p}$ and $V^{\top}_{p}$ into $A$ and $B$ by $A=U_{p}\text{diag}(S_{p})^{\frac{1}{2}}$ and $B=\text{diag}(S_{p})^{\frac{1}{2}}V^{\top}_{p}$, SORSA remains $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ in separate matrices.
+SORSA is defined by Equation 7, which is initially equivalent to the pre-trained weight $W_{0}$.
+During training, $W_{r}$ remains frozen, and only $U_{p}$, $S_{p}$, and $V^{\top}_{p}$ are updated.
+SORSA is defined as:
+$\text{SORSA}(x):=x(W_{r}+W_{p})=xW_{r}+xU_{p}\text{diag}(S_{p})V^{\top}_{p}$ (7)
+In our implementation, we use a optimized version of SORSA equation which results a much faster computation speed, and could approach to the speed of LoRA. See Appendix A for more details.
 We adopt an orthonormal regularizer similar to (Zhang et al., 2023) for $U_{p}$ and $V^{\top}_{p}$. We verify its importance and effectiveness in Section 5.
@@
 The regularizer could keep their orthonormality during training.
-Therefore, parameter updating of $U_{p}$, $\Sigma_{p}$ and $V^{\top}_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
+Therefore, parameter updating of $W_{p}$ in a SORSA adapter at training step $t$ could be expressed as:
 $W_{p,t+1}=$ $W_{p,t}-\eta_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{train}-\gamma_{t}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}$ (9)
@@
 $\eta_{t}$ and $\gamma_{t}$ are the learning rate for training loss and regularizer loss at step $t$, respectively.
-We update the SORSA as Eq equation 10 for implementation simplicity.
+We update the SORSA as Equation 10 for implementation simplicity.
 $W_{p,t+1}=$ $W_{p,t}-\eta_{t}\left(\nabla_{W_{p,t}}{\mathcal{L}}_{train}+\frac{\gamma}{\eta_{d}}\nabla_{W_{p,t}}{\mathcal{L}}_{reg}\right)$ (10)
@@
 We explore four critical aspects: the convexity of the regularizer, the Lipschitz continuity of the gradient, bounds on the hyperparameter $\gamma$, and the impact on the condition number of the optimization problem.
-The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix B.
+The proofs of the theorems and lemmas, along with additional mathematical details, are provided in Appendix C.
 Our analysis reveals fundamental theoretical properties of SORSA, establishing its mathematical soundness and demonstrating its optimization advantages.
 We prove two key theorems that form the cornerstone of our theoretical framework:
-Theorem 5.1.
+Theorem 1.
 The regularizer ${\mathcal{L}}_{reg}$ is convex.
-Theorem 5.2.
+Theorem 2.
 The gradient of the regularizer ${\mathcal{L}}_{reg}$ is Lipschitz continuous.
@@
 Building upon these foundational results, we further analyze the bounds of the hyperparameter $\gamma$, a critical factor in the performance of SORSA:
-Theorem 5.3.
+Theorem 3.
 For convergence of gradient descent, the learning rate $\eta_{d}$ and regularization parameter $\gamma$ must satisfy:
-$\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
+$0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (11)
 This theorem provides crucial guidance for practitioners, offering a clear criterion for selecting appropriate values of $\gamma$ to ensure convergence of the gradient descent process.
@@
 We begin this analysis by establishing a key lemma that bounds the effect of the orthonormal regularizer on the singular values of the weight matrix:
-Lemma 5.4.
-Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\Sigma_{p}^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\Sigma_{p}^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer.
+Lemma 4.
+Let $W_{p}^{\text{unreg}}=U_{p}^{\text{unreg}}\text{diag}(S_{p})^{\text{unreg}}(V_{p}^{\text{unreg}})^{\top}$ be the $W_{p}$ only training without using regularizer, and $W_{p}^{\text{reg}}=U_{p}^{\text{reg}}\text{diag}(S_{p})^{\text{reg}}(V_{p}^{\text{reg}})^{\top}$ be the $W_{p}$ training with the regularizer.
 For each singular value $\sigma_{i}$, the following bound holds:
@@
 where $\epsilon$ is a small positive constant.
-Lemma 5.4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
-Theorem 5.5.
+Lemma 4 provides a crucial connection between the regularizer and the singular values. Building on this result, we arrive at our main theorem regarding the condition number:
+Theorem 5.
 The orthonormal regularizer in SORSA can improve the condition number of the optimization problem over the course of training under certain conditions. Specifically:
@@
 $\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=1$ (13)
-When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$.
-When $\kappa(U_{p,\text{unreg}})\cdot\kappa(V_{p,\text{unreg}}^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
+When training ($t>0$), while $U_{p}^{\text{unreg}}$ and $(V_{p}^{\text{unreg}})^{\top}$ are more and more deviating from orthonormal, $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ will increase, and $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$ will generally remain $1$ due to the regularization in orthonormality of $U_{p}^{\text{reg}}$ and $(V_{p}^{\text{reg}})^{\top}$.
+When $\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})$ becomes greater than $(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})$, the ratio becomes strictly less than 1, indicating an improvement in the condition number.
 This theorem quantifies the improvement in the condition number achieved by SORSA, offering an explanation for its fast convergence.
@@
 This theorem could also shows that training with the regularizer, the distribution of $W_{p}$ will be more evenly distributed due to a smaller ratio between $\sigma_{\max}(W_{p})$ and $\sigma_{\min}(W_{p})$, which means a better training stability.
+Moreover, as mentioned in Büyükakyüz (2024), orthonormal matrices in neuron networks could improve gradient flow (Saxe et al., 2014; Arjovsky et al., 2016) and enhanced optimization landscape (Huang et al., 2018; Wisdom et al., 2016), which could also explain SORSA’s superior performance in convergence.
 In conclusion, these theorems provide a mathematical foundation for the SORSA method.
@@
 We conduct comparative experiments on different NLP tasks, including natural language generation (NLG) between SORSA, PiSSA (Meng et al., 2024), LoRA (Hu et al., 2021), and full parameter fine-tuning.
-We conducted NLG tests on Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023).
-We trained the models with SORSA applied using MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
-We also trained the model on CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021).
+We conducted NLG tests on Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024).
+We trained the models using first 100K data in MetaMathQA (Yu et al., 2024), and evaluated the model on GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021).
+We also trained the model on first 100K data in CodeFeedback Filtered Instruction (Zheng et al., 2024) dataset and evaluated it on HumanEval (Chen et al., 2021).
 The training process followed identical setups as the experiments conducted in PiSSA (Meng et al., 2024).
 All reported values are accuracy.
-See Appendix A.2 for more details and hyperparameters of the training.
-We quoted results for PiSSA, LoRA, and full parameter fine-tuning directly from Meng et al.
+See  Section B.2 for more details and hyperparameters of the training.
+We quoted some results of PiSSA, LoRA, and full parameter fine-tuning from Meng et al.
 (2024).
-For SORSA experiments, we conducted our experiments on an NVIDIA A100-SXM4 (80GB) GPU.
+For experiments, we conducted part of our experiments on a single NVIDIA A100-SXM4 (80GB) GPU, others on a single NVIDIA H100-SXM4 (80GB) GPU.
 See Table 1 for the results and Figure 3 for the loss and gradient norm comparison.
-Model Method Trainable Parameters MATH GSM-8K HumanEval
-Llama 2 7B Full FT 6738M 7.22 49.05 21.34
-LoRA 320M 5.50 42.30 18.29
-PiSSA 320M 7.44 53.07 21.95
-SORSA 320M 10.36 56.03 24.39
-Mistral 7B Full FT 7242M 18.60 67.02 45.12
-LoRA 168M 19.68 67.70 43.90
-PiSSA 168M 21.54 72.86 46.95
-SORSA 168M 21.86 73.09 47.56
+Model Method Trainable Parameters GSM-8K MATH HumanEval
+Llama 2 7B Full FT 6738M 49.05 7.22 21.34
+LoRA 320M 42.30 5.50 18.29
+PiSSA 320M 53.07 7.44 21.95
+SORSA 320M 56.03 10.36 24.39
+RWKV6 7B LoRA 176M 8.04 7.38 15.24
+PiSSA 176M 32.07 9.42 17.07
+SORSA 176M 45.87 11.32 22.56
+Mistral 7B Full FT 7242M 67.02 18.60 45.12
+LoRA 168M 67.70 19.68 43.90
+PiSSA 168M 72.86 21.54 46.95
+SORSA 168M 73.09 21.86 47.56
+Gemma 7B Full FT 8538M 71.34 22.74 46.95
+LoRA 200M 74.90 31.28 53.66
+PiSSA 200M 77.94 31.94 54.27
+SORSA 200M 78.09 29.52 55.49
 Table 1: Comparing SORSA with other methods on NLG tasks. denotes results from Meng et al. (2024).
 Refer to caption Refer to caption
-Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of Llama 2 7B. LoRA and PiSSA curves are from Meng et al. (2024).
-The results showed that on the Llama 2 7B model, SORSA achieved 10.36% accuracy on the MATH benchmark and 56.03% accuracy on the GSM-8K benchmark, both significantly outperforming PiSSA, LoRA, and Full FT.
-On the Mistral 7B v0.1 model, SORSA achieved 21.86% accuracy on the MATH benchmark and 78.03% accuracy on the GSM-8K benchmark, demonstrating slightly better performance on both evaluation tasks compared to the other three methods.
-In this section, due to the limitation of computing resources, we train and benchmark a small number of tasks.
+Refer to caption Refer to caption
+Figure 3: The training loss and gradient norm comparison between SORSA, PiSSA, and LoRA on MetaMathQA training of RWKV6 7B and Llama 2 7B.
+LoRA and PiSSA curves of Llama 2 7B are from Meng et al.
+(2024).
+The results showed that across all models tested, SORSA generally outperformed other methods.
+For mathematical evaluations, on Llama 2 7B, SORSA scored 56.03% on GSM-8K and 10.36% on MATH, significantly outperforming other methods; For the RWKV6 7B model, SORSA achieved 45.87% accuracy on GSM-8K and 11.32% on MATH, surpassing both PiSSA and LoRA; On Mistral 7B, SORSA reached 73.09% on GSM-8K and 21.86% on MATH, slightly edging out competitors; With Gemma 7B, while SORSA’s MATH score of 29.52% was lower than LoRA and PiSSA, it still led on GSM-8K with 78.09%.
+Moreover, on coding evaluations, SORSA consistently achieved the highest HumanEval scores across other tuning methods, demonstrating its effectiveness on diverse natural language tasks.
+Noticeably, RWKV6 7B using LoRA performed extremely below our expectation in GSM-8K test.
+We believe this may be due to our prompt for mathematical tests (See Section B.2 for more details) are inconsistent with the model’s original prompt, causing LoRA to fail in learning this prompting behavior.
+The Figure 3 reveals that while SORSA and PiSSA exhibit nearly identical loss curves at the beginning, and even a little bit higher than PiSSA on RWKV-6 training.
+However, when training step approximately $t>300$, SORSA continues to maintain the decrease of its loss steadily.
+In contrast, LoRA and PiSSA shows a deceleration in its loss reduction.
+This supports Theorem 5, confirming that SORSA consistently outperforms both LoRA and PiSSA, especially at later stages of training.
+However, due to the limitation of computing resources, we only trained and benchmarked a small number of tasks.
 7 Conclusion
@@
 Specifically, Llama 2 7B, tuned with SORSA, achieved significant improvements in the GSM-8K and MATH benchmarks, highlighting the effectiveness of our approach.
-We adopt singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training.
+We adopted singular values and vectors analysis, comparing SORSA with partial FT and LoRA, and we find SORSA’s superiority in preserving the pre-trained weight’s singular values and vectors during training.
 This suggests an explanation for SORSA’s supreme performance demonstrated in the experiment.
 We also show the significance of the orthonormal regularizer through analysis.
-Our gradient analysis provides a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape.
+Our gradient analysis provided a mathematical foundation for SORSA, demonstrating its convexity, Lipschitz continuity, and the crucial role of the regularizer in improving the optimization landscape.
 This theoretical framework not only explains SORSA’s empirical superior performance but also offers valuable insights for future developments in adaptive learning algorithms.
@@
 While SORSA demonstrates substantial improvements over existing PEFT methods, there are several avenues for future work to enhance its capabilities further and extend its applicability:
-  • Extended Evaluation: Conduct a broader evaluation of SORSA across various benchmarks in different models, in order to provide a more comprehensive understanding of SORSA’s performance and versatility.
+  • Comparison Between Different Rank: Find how the performance of SORSA varies with different ranks.
+This investigation would help determine the optimal rank for various tasks and model sizes, potentially revealing a trade-off between computational efficiency and model performance.
+  • Extended Evaluation: Conduct broader evaluations of SORSA across different areas of benchmarks on models with more diverse sizes.
+This may able to provide a more comprehensive understanding of SORSA’s performance and versatility.
   • Catastrophic Forgetting: Compare the performance of some general benchmarks like MMLU (Hendrycks et al., 2020) before and after models are adapted by SORSA in specific down-stream tasks in order to research whether SORSA could perform better than full-parameter fine-tuning, LoRA (Hu et al., 2021) or PiSSA (Meng et al., 2024) on preventing catastrophic forgetting.
-  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024).
-This research aims to reduce the memory footprint and computational requirements of SORSA, making it feasible to deploy on edge devices or in environments with limited computational resources.
-By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, QSORSA could enable highly efficient and powerful language models suitable for a broader range of practical applications.
+  • Quantization: Investigate the application of quantization techniques to SORSA similar to the approach of QLoRA (Dettmers et al., 2024) and QPiSSA (Meng et al., 2024).
+Quantization aims to reduce the memory footprint and computational requirements, making it feasible to deploy on devices with limited computational resources.
+By combining the benefits of quantization with the efficient fine-tuning capabilities of SORSA, quantization of SORSA may able to efficiently provides powerful models, which could eventually facilitate AI for a broader range of practical applications.
 By pursuing these directions, future work can build on the foundation laid by SORSA, pushing the boundaries of PEFT and enhancing the adaptability and performance of large language models across a wide range of applications.
@@
 9 Impact Statement
-In this paper, we introduce an innovative PEFT method in the area of Machine Learning.
-Our approach significantly streamlines the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability.
+In this paper, we introduced an innovative PEFT method in the area of Machine Learning.
+Our approach significantly streamlined the model’s tuning process, particularly for large-scale models, addressing both computational efficiency and environmental sustainability.
 As we push the boundaries of what is possible with Machine Learning, it is essential to consider the broader impacts of these advancements on both the environment and ethical standards within the field.
 9.1 Environmental Impact
-SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B model than full parameter fine-tuning.
-This significant reduction in hardware resource requirements translates to approximately 50% less energy consumption compared to full parameter fine-tuning methods.
-By enhancing efficiency, our approach contributes to reducing the environmental footprint of machine learning operations.
+SORSA optimizes model tuning for large-scale models, reducing VRAM consumption by over 50% when adapting the Llama 3 70B Dubey et al.
+(2024) model than full parameter fine-tuning.
+This significant reduction in hardware resource requirements also suggests a less energy consumption compared to full parameter fine-tuning methods.
+By enhancing efficiency, our approach could significantly contributes to reducing the carbon footprint of Machine Learning operations.
 9.2 Ethical Concerns
@@
 Jianyong Wang from Tsinghua University and the Computer Science Department of the Advanced Project Research Laboratory of Tsinghua University High School for providing the endorsement to the research.
-Appendix A Experiments Details
-A.1 Analysis
+Appendix A Faster SORSA Adapters
+According to the definition of SORSA from Equation 7, because $\text{diag}(S_{p})$ is always a diagonal matrix, it is equivalent to:
+$\text{SORSA}(x)=xW_{r}+x(U_{p}\odot S_{p})V_{p}^{\top}$ (14)
+Where $\odot$ denotes element-wise multiplication, and $\text{diag}(W)$ denotes the diagonal elements of the matrix $W$.
+This transformation allows us to reduce the computational complexity of SORSA adapters.
+In the original form, we have to perform twice matrix multiplications.
+However, in the Equation 14, we only have one matrix multiplication, and one element-wise multiplication.
+The time complexity of element-wise multiplication is $O(m^{2})$, which is much less than matrix multiplication’s complexity $O(m^{3})$.
+Therefore, while $\lim_{m\to\infty}$, the computation speed of SORSA adapters will be same as LoRA and PiSSA.
+We performed a benchmark using PyTorch (Paszke et al., 2019) on a NVIDIA H100 SXM4 (80GB) GPU backed with CUDA, and Apple M2Pro CPU, in order to test the computation time between these two methods.
+See Figure 4 to see our results.
+Refer to caption Refer to caption
+Figure 4: Benchmark between two equations of SORSA
+Appendix B Experiments Details
+B.1 Analysis
 For the singular values and vectors analysis in Section 3, we applied partial fine-tuning, LoRA and SORSA (with and without orthonormal regularizer) on Llama 2 7B (Touvron et al., 2023) model, training with the first 100K data in MetaMathQA (Yu et al., 2024) dataset.
+We only calculated the loss on the response part.
+The models are trained with TF32 & BF16 (Wang & Kanwar, 2019) mix precision.
+See Table 2 for hyperparameters.
+We used AdamW (Loshchilov & Hutter, 2018) optimizer and cosine annealing scheduler in training.
 In the analysis, LoRA and SORSA were only applied to q_proj and v_proj matrices, respectively.
 For Partial FT, we set model’s q_proj and v_proj matrices to trainable.
-We only calculated the loss on the response part.
-The models are trained with TF32 + BF16 mix precision.
-See Table 2 for hyperparameters.
 We also found we should only perform SVD for analysis using CPU, in order to get the precise analysis data.
+Model Llama 2 7B
+Method Partial FT LoRA SORSA (w/o reg) SORSA
 Training
-Method FT LoRA SORSA (w/o reg) SORSA
-Precision TF32 + BF16 Mix-Precision
-Epoch 1
-Batch Size 128
-Max Length 512
-Weight Decay 0.00
-Warm-up Ratio 0.03
+Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
+Epoch 1 1 1 1
+Batch Size 128 128 128 128
+Max Length 512 512 512 512
+Weight Decay 0 0 0 0
+Warm-up Ratio 0.03 0.03 0.03 0.03
 Learning Rate 2e-5 2e-5 2e-5 3e-5
-$\gamma$ 0 5e-4
-Rank 128 128 128
+Grad Clip False False False False
+SORSA $\gamma$ N/A N/A 0 5e-4
+Rank N/A 128 128 128
 Table 2: Hyperparameters for the analysis
-A.2 NLG Experiments
-For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023) and Mistral 7B v0.1 (Jiang et al., 2023) models by SORSA.
+B.2 NLG Experiments
+For our NLG tasks, we adapted Llama 2 7B (Touvron et al., 2023), RWKV6 7B (Peng et al., 2024), Mistral 7B v0.1 (Jiang et al., 2023) Gemma 7B (Gemma Team et al., 2024).
+models by SORSA.
 For GSM-8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) evaluations, we trained those models with the first 100K data in MetaMathQA (Yu et al., 2024) dataset.
@@
 We only calculated the loss on the response part.
-The models are trained with TF32 + BF16 mix precision.
+The models are loaded in FP32, and trained with TF32 & BF16 mix precision.
+In our experiments, we selected a higher learning rate for SORSA compared to other methods to counterbalance the negative effect of orthonormal regularizer on optimizing toward lower training loss.
 See Table 3 for hyperparameters.
+See LABEL:lst:math for the prompt we used in GSM-8K and MATH evaluations, and LABEL:lst:humaneval for the prompt we used for HumanEval tests.
+Model Llama 2 7B Mistral 7B Gemma 7B RWKV6 7B RWKV6 7B
+Method SORSA SORSA SORSA SORSA LoRA PiSSA
 Training
-Model Llama 2 7B Mistral 7B v0.1
-Method SORSA
-Precision TF32 + BF16 Mix-Precision
-Epoch 1
-Batch Size 128
-Max Length 512
-Weight Decay 0.00
-Warm-up Ratio 0.03
-Learning Rate 3e-5
-$\gamma$ 4e-4
-Rank 128 64
-Inference
-Precision BF16
+Mix-Precision TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16 TF32&BF16
+Epoch 1 1 1 1 1
+Batch Size 128 128 128 128 128
+Max Length 512 512 512 512 512
+Weight Decay 0 0 0 0 0
+Warm-up Ratio 0.03 0.03 0.03 0.03 0.03
+Learning Rate 3e-5 3e-5 3e-5 3e-5 2e-5
+Grad Clip 1.0 1.0 1.0 1.0 1.0
+SORSA $\gamma$ 4e-4 4e-4 4e-4 4e-4 N/A
+Rank 128 64 64 64 64
+Evaluating
+Precision BF16 BF16 BF16 FP32 FP32
+Sampling False
 Top-P 1.0
-GSM-8K Max Length 1024
-MATH Max Length 2048
-HumanEval Max Length 2048
+Max Length GSM-8K: 1024 MATH: 2048 HumanEval: 2048
 Table 3: Hyperparameters of training on models for GSM-8K and MATH
+1Below is an instruction that describes a task. Write a response that appropriately completes the request.
+2
+3### Instruction:
+4{question}
+5
+6### Response: Lets think step by step.
+Listing 1: Prompt used for GSM-8K and MATH.
 1@@ Instruction
@@
 3‘‘‘python
-4{prompt}
+4{question}
 5‘‘‘
@@
 15@@ Response
-Listing 1: Prompt used for HumanEval evaluation.
-Appendix B Proofs
-B.1 Proof of Theorem 5.1
+Listing 2: Prompt used for HumanEval evaluation.
+Appendix C Proofs
+See 1
 Proof.
@@
 Therefore, both $f(U_{p})$ and $g(V_{p})$ are convex, and consequently, ${\mathcal{L}}_{reg}$ is convex. ∎
-B.2 Proof of Theorem 5.2
+See 2
 Proof.
@@
 This proves that ${\mathcal{L}}_{reg}$ is Lipschitz continuous with Lipschitz constant $L_{reg}$. ∎
-B.3 Proof of Theorem 5.3
+See 3
 Proof.
@@
 Conversely, smaller steps lead to a more gradual progression towards orthonormality, which reduces $L_{reg}$.
-Therefore, we could assume $L_{reg}\propto\eta_{d}$, and rewrite the inequality as:
+Therefore, we could assume $L_{reg}\propto\eta_{d}$.
+Moreover, the $\gamma$ must not be negative or the regularization term would impact negatively on the purposes it is supposed to have.
+Therefore, we can rewrite the inequality as:
 $0\leq\gamma<\frac{2}{k\eta_{d}}-L_{train}$ (23)
 This establishes the upper bound for $\gamma$. ∎
-B.4 Proof of Lemma 5.4
+See 4
 Proof.
 First, let’s consider the effect of the orthonormal regularizer. The regularizer aims to make $U_{p}^{\top}U_{p}\approx I$ and $V_{p}V_{p}^{\top}\approx I$. We can quantify this approximation as:
-$\nabla_{W_{p}}{\mathcal{L}}_{reg}\leq\epsilon_{\nabla}$ (23)
+$\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}\leq\epsilon_{\nabla}$ (24)
 where $\epsilon_{\nabla}>0$ is a small constant.
 Now, let’s consider the difference between $W_{p}^{\text{unreg}}$ and $W_{p}^{\text{reg}}$:
-$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}$ (24)
-Now, we can use Weyl’s inequality, which states that for matrices A and B:
+$\|W_{p}^{\text{reg}}-W_{p}^{\text{unreg}}\|_{F}=\gamma\|\nabla_{W_{p}}{\mathcal{L}}_{reg}\|_{F}=\gamma\epsilon_{\nabla}$ (25)
+Now, we can use Weyl’s inequality (Weyl, 1912), which states that for matrices A and B:
 $|\sigma_{i}(A+B)-\sigma_{i}(A)|\leq\|B\|_{2}\leq\|B\|_{F}$ (26)
@@
 $(1-\epsilon)\sigma_{i}^{\text{unreg}}\leq\sigma_{i}^{\text{reg}}\leq(1+\epsilon)\sigma_{i}^{\text{unreg}}$ (29)
-B.5 Proof of Theorem 5.5
+See 5
 Proof.
-Let $W_{p}=U_{p}\Sigma_{p}V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
+Let $W_{p}=U_{p}\text{diag}(S_{p})V_{p}^{\top}$ be the principal part of the singular value decomposition approximation of $W$. The condition number is given by:
 $\kappa(W_{p})=\frac{\sigma_{\max}(W_{p})}{\sigma_{\min}(W_{p})}$ (30)
@@
 And $\epsilon_{0}=0$, $\delta_{1,0}=\delta_{2,0}=0$. Therefore:
-$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\Sigma_{p}^{\text{reg}})}{\kappa(\Sigma_{p}^{\text{unreg}})}=1$ (31)
+$\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}=\frac{\kappa(\text{diag}(S_{p})^{\text{reg}})}{\kappa(\text{diag}(S_{p})^{\text{unreg}})}=1$ (32)
 During training ($t>0$): As training progresses, $U_{p}$ and $V_{p}^{\top}$ deviate from orthonormality in the unregularized case. We quantify this deviation:
@@
 where $\delta_{1,t},\delta_{2,t}$ are small positive numbers that remain bounded due to the regularization.
-From the Lemma 5.4, we arrive at:
+From the Lemma 4, we arrive at:
 $\begin{split}\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1-\epsilon_{t}}{1+\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\leq\frac{\kappa(W_{p}^{\text{reg}})}{\kappa(W_{p}^{\text{unreg}})}\leq\frac{(1+\delta_{1,t})(1+\delta_{2,t})(\frac{1+\epsilon_{t}}{1-\epsilon_{t}})}{\kappa(U_{p}^{\text{unreg}})\cdot\kappa((V_{p}^{\text{unreg}})^{\top})}\end{split}$ (37)